### v2

# AUTALIC: A Dataset for Anti-AUTistic Ableist Language In Context

Abstract
As our understanding of autism and ableism continues to increase, so does our understanding of ableist language towards autistic people. Such language poses a significant challenge in NLP research due to its subtle and context-dependent nature. Yet, detecting anti-autistic ableist language remains underexplored, with existing NLP tools often failing to capture its nuanced expressions. We present AUTALIC, the first benchmark dataset dedicated to the detection of anti-autistic ableist language in context, addressing a significant gap in the field. The dataset comprises 2,400 autism-related sentences collected from Reddit, accompanied by surrounding context, and is annotated by trained experts with backgrounds in neurodiversity. Our comprehensive evaluation reveals that current language models, including state-of-the-art LLMs, struggle to reliably identify anti-autistic ableism and align with human judgments, underscoring their limitations in this domain. We publicly release AUTALIC along with the individual annotations which serve as a valuable resource to researchers working on ableism, neurodiversity, and also studying disagreements in annotation tasks. This dataset serves as a crucial step towards developing more inclusive and context-aware NLP systems that better reflect diverse perspectives.
AUTALIC: A Dataset for Anti-AUTistic Ableist Language In Context
Naba Rizvi1, Harper Strickland1, Daniel Gitelman1, Tristan Cooper1, Alexis Morales-Flores1, Michael Golden1, Aekta Kallepalli1, Akshat Alurkar1, Haaset Owens1, Saleha Ahmedi1, Isha Khirwadkar1, Imani Munyaka1, Nedjma Ousidhoum2 1 University of California, San Diego, 2 Cardiff University, Correspondence: nrizvi@ucsd.edu
Figure 1: The example illustrates the importance of labeling sentences in context. The target sentence alone, shown on the left, is difficult to classify as ableist toward autistic people. Adding the surrounding sentences, as shown on the right, provides context revealing the original poster’s reference to the debunked vaccine-autism stereotype, which is tied to anti-autistic stigma Mann (2019); Davidson (2017).
Trigger warning: this paper contains examples of anti-autistic ableist language including explicit slurs and references to violence.
1 Introduction
There are several critical frameworks used to define and understand disability, including autism Lawson and Beckett (2021). While these frameworks are ever-evolving, the medical model framework which defines disability as a “disease” is one of the most commonly applied in computer science research when discussing disability in general and autism specifically Rizvi et al. (2024); Spiel et al. (2019a); Sideraki and Drigas (2021); Anagnostopoulou et al. (2020); Parsons et al. (2020); Williams et al. (2023); Sum et al. (2022). This framework defines autism as a deficit of skills such as the Theory of Mind, and its applications in technology research largely focus on providing diagnosis and treatment to autistic people Baron-Cohen (1997); Begum et al. (2016); Rizvi et al. (2024); Spiel et al. (2019b). However, this framework has been criticized for creating a power imbalance between medical professionals and the autistic community as it promotes the belief that a formal diagnosis is necessary to validate the autistic identity Bennett and Keyes (2020). This belief also posits neurotypical behaviors as the “norm” and autism as a “deficit” of these norms, thereby promoting neuronormativity instead of neurodiversity, which views all neurotypes as valid forms of human diversity Bottema-Beutel et al. (2021); Walker (2014).
In contrast, prior work has shown viewing autism as an identity may increase the psychological well-being of autistic individuals and lower their social anxiety Cooper et al. (2023). Although the language used to describe autism varies, prior studies with autistic American adults found 87% prefer identity-first language over person-first language Taboas et al. (2023). Person-First Language (PFL) centers the person (i.e. person with autism), while Identity-First Language (IFL) centers the identity (i.e. autistic person) Taboas et al. (2023).
Prior work examining ableist speech directed at content creators identified 11 types of “ableist hate and harassment" that they categorized as slurs and derogatory language, violent and eugenics-related speech, questioning ability and denying access, mocking and invalidation of disability identity, and objectifying the disabled body Heung et al. (2024). Additionally, researchers have identified emerging terms such as “weaponized autism" that are gaining popularity on alt-right social media platforms, further reflecting the ever-evolving nature of negative societal attitudes towards autism Welch et al. (2023).
Anti-autistic ableist language can be diverse in scope. It may include perpetuating stereotypes, using offensive language and slurs, or centering non-autistic people over the perspectives of autistic people Bottema-Beutel et al. (2021); Rizvi et al. (2024); Darazsdi and Bialka (2023). While abusive language detection systems can help identify such speech, they are known to demonstrate bias Manerba and Tonelli (2021); Venkit et al. (2022), with even LLMs perpetuating ableist biases Gadiraju et al. (2023a). Additionally, anti-autistic ableist speech remains understudied, which is concerning given that classifiers trained on multiple hate speech datasets have shown a failure to generalize to target groups outside of the training corpus Yoder et al. (2022).
Additionally, no datasets specifically focus on anti-autistic speech classification, and only 3 of the 23 datasets for bias evaluation in LLMs focus on disability Gallegos et al. (2024). Prior work has also found that LLMs lack an acknowledgment of context, which leads to higher rates of false positives when classifying ableist speech Phutane et al. (2024). These limitations are also found in toxicity classifiers, which excel primarily at identifying explicit ableist speech but may otherwise perpetuate harmful social biases leading to content suppression Phutane et al. (2024).
Our work addresses these gaps in existing research by examining the performance of our dataset in the following domains for the task of anti-autistic hate speech classification:
  1. 1.
    The impact our dataset has on fine-tuning classifiers
  2. 2.
    The alignment of LLMs with the perspectives of trained human annotators
  3. 3.
    The extent to which in-context learning improves human-LLM alignment.
2 Related Work
Prior studies have also collected insights on anti-autistic language from the parents of autistic children Allouch et al. (2018, 2019). However, toxic language datasets focusing on hate speech and abusive language have often addressed disability in general terms but have not explicitly focused on autism ElSherief et al. (2018); Ousidhoum et al. (2019). Therefore, toxic language detection models, including LLM-based models, have been found to exhibit strong negative biases toward disabilities by classifying any disability-related text as toxic Narayanan Venkit et al. (2023). Further, LLMs have been observed to perpetuate implicitly ableist stereotypes Gadiraju et al. (2023b) and bias Gama (2024); Venkit et al. (2022). This, unfortunately, can sometimes be due to a research design that overlooks intra-community and disabled people’s perspectives Mondal et al. (2022), as well as autistic people’s views, which may lead to harmful stereotypes Rizvi et al. (2024); Spiel et al. (2019b). We make a step towards addressing these issues by building a dataset that focuses on ableist speech and autism by including autistic people’s perspectives during the annotation processing as recommended by Davani et al. (2023). AUTALIC contains all its labels and will also be useful for researchers interested in leveraging disagreements for difficult classification tasks Leonardelli et al. (2021); Pavlick and Kwiatkowski (2019).
3 AUTALIC
This section discusses the process of compiling and annotating the AUTALIC dataset. We collected relevant sentences containing autism-related keywords from Reddit using the methods described in Section 3.1. The collected sentences were labeled by trained annotators, as discussed in Section 3.2.
3.1 Data Collection
Here, we present details from our data collection process, which include identifying our methodology for gathering representative sentences related to autism, curating the data obtained through our search parameters, and summarizing the resulting AUTALIC dataset.
3.1.1 Data Collection Criteria
First, we identified potential data sources and selected Reddit as the preferred source based on its popularity, focus on text-based content, and fewer API restrictions than X at the time of our data collection. We searched for keywords using the default search settings on Reddit, which filters posts based on relevancy by prioritizing rare words in the search query, the age of the post, and the amount of likes and comments it has   [^1] https://support.reddithelp.com/hc/en-us/articles/19695706914196-What-filters-and-sorts-are-available. Table 1 displays the keywords we searched for on Reddit and the final number of sentences with each keyword in our dataset.
Word Sentence Count
autis* 1,221
ASD 226
disabilit* 184
aspergers 173
ABA 167
neurotyp* 158
aspie* 144
neurodiver* 103
AuDHD 99
disable* 93
autism speaks 66
stupid* 56
a11y 34
NT 27
retard* 25
idiot* 18
actually autistic 13
autism intensifies 6
ND 5
autie* 2
Table 1: The keywords included in our Reddit search and the number of sentences containing each term in the AUTALIC dataset.
We use the identified search terms detailed in Table  1 to collect two types of sentences:
  • the target instance containing our keywords, to be labeled by the annotators,
  • the sentences preceding or following target instances to provide additional context.
We collected 2,500 target sentences, with 2,014 preceding and 2,500 following them. Finally, we split our dataset into three parts by randomly selecting and assigning 800 unique target sentences to create three segments that were each annotated by a group of three annotators.
3.1.2 Data Curation
As some of the identified keywords may appear in other contexts, we perform an exact word search for the acronyms to ensure unrelated words that might contain our acronyms are excluded from the search as they go beyond the scope of our dataset. For example, we searched for “applied behavioral analysis” and a case-sensitive search for ABA, which is a form of therapy intended to minimize autistic behaviors such as stimming (which is often used for self-soothing) Sandoval-Norton et al. (2019). Similarly, we exclude any posts that are not written in English using the Python package langdetect  [^2] https://pypi.org/project/langdetect/ and posts that contain images, videos, or links.
3.1.3 Final Dataset
Our final dataset includes 2,400 sentences from 192 different subreddits. To protect our annotators’ privacy, we have anonymized all labels.
While nearly a quarter of the posts in our dataset were published in 2023, the range of publication years is 2013-2024. Figure 2 shows an example of a sentence from our dataset that uses both a search keyword and a word defined in our glossary described in Section 3.2.2.
Figure 2: An example of a sentence from our dataset. The search keyword is shown in red, while the word in blue is an example of a word defined in our glossary.
The average number of likes on each post included in the AUTALIC dataset is 1,611.59, and only 135 posts have 0-5 likes. Table 2 details the subreddits from which we extracted the most significant number of sentences for our dataset.
SubReddit Sentence Count
r/Aspergers 116
r/Autism 88
r/AmITheAsshole 39
r/AutisminWomen 37
r/AuDHDWomen 24
Table 2: The subreddits with the most sentences included in the AUTALIC dataset and the number of sentences extracted from each.
3.2 Data Annotation
After obtaining sample sentences with contextual surrounding sentences, we proceeded to annotation: annotator selection, training, and data annotation.
3.2.1 Annotator Selection
We recruited nine annotators and randomly assigned them to annotate different segments of the dataset. Our annotators are undergraduate students who volunteered to annotate the data without financial compensation and were offered authorship following the ACL guidelines on authorship. They are culturally diverse and include people who grew up outside the United States. However, they are all fluent in English. Four of our annotators are gender minorities, and at least three self-identify as neurodivergent.
Although we ensured the annotators were from diverse backgrounds during our recruitment process, due to the collaborative nature of our annotation process, we cannot share the individual details of their identities. We also note that any personally identifiable information was destroyed upon the conclusion of our analysis and not shared outside of our research team.
3.2.2 Annotator Training
We provided a virtual orientation to all annotators explaining the history of anti-autistic ableism, examples of contemporary anti-autistic discrimination, and a brief overview of the annotation task.
The orientation begins with a discussion of the medical model approach to autism and its link to the Nazi eugenics program Waltz (2008); Sheffer (2018). We define neuronormativity as the belief that the neurotypical brain is “normal” and other neurotypes are deficient in neurotypicality Wise (2023). We dive deeper into the medical model by discussing its impact on the self-perceptions and inclusion of autistic people in our society, such as an increase in suicidal ideation and social isolation among autistic people who mask or hide their autistic traits  Cassidy et al. (2014, 2018). Then, we cover the shifts in perspectives that emerged due to disability rights activism  Rowland (2015); Cutler (2019), and define neurodiversity as the belief that all neurotypes are valid forms of human diversity  Walker (2014).
To explain the annotation task, we provide examples of sentences similar to what they may encounter while annotating. For example, we discuss how the inclusion of "at least" alters the connotations of the follow sentence:
At least I am not autistic.
With just a minor change, the sentence can have an ableist connotation as it implies relief in knowing one is not autistic, as if it is shameful or wrong.
We also introduce our glossary to the annotators as a dynamic resource that can be altered as needed. This glossary contains words that may appear in autism discourse online that may not be commonly known to others. These include medical acronyms, slang, and references to organizations and resources commonly affiliated with the autistic community (such as Autism Speaks). Selected examples from our glossary are shown in Table  3. The final version of our glossary is available can be found in this footnote   [^3] https://figshare.com/s/3850e63295ccf6fceebb .
Term Definition
AuDHD A combination of ADHD and autism
ND Neurodivergent or neurodiverse
NT Neurotypical
Table 3: Selected examples from our glossary detailing commonly used acronyms, abbreviations, and slang in autism discourse online. For the full glossary, see footnote 3.
We conclude our orientation by providing a brief tutorial video demonstrating how to run the script that will guide each annotator through the annotation task.
3.2.3 Data Labeling
After completing the training, we assign each of the three segments of the dataset to three randomly selected annotators. Each annotator is assigned 800 unique sentences, with a goal of completing 200 annotations each week over four weeks.
Annotators select from three possible labels for each sentence: “Ableist,” “Not Ableist,” or “Needs More Context.” According to the Center for Disability Rights, “Ableism is a set of beliefs or practices that devalue and discriminate against people with physical, intellectual, or psychiatric disabilities and often rests on the assumption that disabled people need to be ‘fixed’ in one form or the other.” [^4] https://cdrnys.org/blog/uncategorized/ableism/ We supplied this definition to annotators and instructed them to select the “ableist” label for any sentences that contain examples of these beliefs and practices directed towards autistic people.
Annotators are instructed to select the “not ableist” label for sentences that describe positive behaviors and attitudes regarding autism or disabilities or may be written by an autistic person reaching out for help and support. This includes individuals using medical terminology in a personal context (e.g., ‘I need therapy’), discussion and suggestions from neurodivergent people (community-generated discussions), and general discussions of the medical processes (unrelated to neurodivergence). Some examples of statements labeled as not ableist are: “I am autistic”, “As an autistic person, I think…”
The “Needs More Context” label is used for sentences an annotator is unable to definitively categorize as ableist or not ableist with only the surrounding sentences as context. This category includes text that is entirely unrelated to disabilities or remains ambiguous without additional context.
The labels used in our annotation task are -1 for sentences that are unrelated/need more context, 0 for sentences that are not anti-autistic, and 1 for sentences that are anti-autistic. The number of times our annotators assigned each label in our dataset is detailed in Table 4. To calculate our agreement scores, we consolidated labels -1 and 0 together based on feedback from our annotators that unrelated sentences needing more context could be classified as not anti-autistic in a purely granular classification.
Label Definition Count
-1 unrelated to autism or needs more context 595
0 not ableist 5,582
1 ableist 1,023
Table 4: An explanation of the labels used in our classification task and the resulting counts of each label from all annotators combined.
While we use the majority label as the ground truth in our analysis, in our public dataset, we will be releasing the individual labels from each annotator due to a growing interest in embracing disagreements for such classification tasks in NLP Leonardelli et al. (2021); Kralj Novak et al. (2022); Pavlick and Kwiatkowski (2019).
3.2.4 Providing Context
While we provide additional sentences for context, the annotators are instructed to annotate the target sentence exclusively and only refer to the other sentences for additional context, such as determining whether the sentence is part of an intra-community discussion or the use of figurative speech (i.e. sarcasm). Figure 1 provides an example of a target sentence in context.
In this example, it is difficult to determine whether or not the writer had ableist intent, as it can be interpreted in multiple ways. For example, they can be critiquing the medical model, as many autistic activists do, thereby making it non-ableist. Or they could be genuinely promoting ableist misrepresentations. The contextual sentences help the annotators better understand the writer’s intent. With these sentences, it is apparent that the writer is referring to the harmful and widely discredited association of vaccines with autism, which not only promotes anti-autistic ableism in society but also puts people’s lives at risk by spreading disinformation about the benefits and harms of life-saving vaccines Gabis et al. (2022); Taylor et al. (2014); Hotez (2021).
Throughout the annotation process, annotators can edit previous annotations based on new knowledge to account for changes in language usage and connotations and the annotators’ dynamic understanding of ableism.
3.2.5 Disagreements
The Fleiss’s Kappa scores were $0.125$ for Team 1, $0.307$ for Team 2, and $0.191$ for team 3. These scores underline the difficulty of our classification task, which is apparent from the findings of prior work Ousidhoum et al. (2019). We analyzed the sentences with the highest levels of disagreement in our dataset. We found that such posts:
  1. 1.
    tend to use medical model terminology or stereotypes
  2. 2.
    need additional context beyond the sentences we provided.
Our dataset contains 59 of these sentences.
Figure 3: An example of a sentence from our dataset, shown in red, with a high level of disagreement among annotators.
Figure 3 contains an example of a sentence with a high disagreement among our annotators. While functioning labels are considered ableist due to their eugenicist approach of categorizing autistic people based on their perceived economic value  De Hooge (2019), it is difficult to determine whether the original poster is autistic or not. The context is important here as classifying a sentence such as this as ’ableist’ can lead to unfair censorship if the original poster is a self-diagnosed autistic person seeking advice. Therefore, these sentences were ultimately classified as ’not ableist’ in AUTALIC.
Our analysis reveals a moderately strong negative correlation between the task completion time and agreement with other annotators that is statistically significant ($R=-0.644$, $p$-value: $0.0096$). This highlights the importance of our orientation as we provided it simultaneously to the annotators, and the annotators who completed their task immediately after our orientation had higher agreement.
4 Experiments
4.1 Data
Our dataset contains 2,400 sentences labeled as containing anti-autistic ableist language or not. The labels are obtained by calculating the mode from the three annotators of each data segment. Using this methodology, 242 target sentences contain examples of anti-autistic ableist language (10% of total), and 2160 sentences do not (90% of total).
4.2 Experimental Setup
We test the performance of different types of models on our dataset, including logistic regression, BERT, and LLMs ranging from PHI 3 Mini to GPT 4. These models are selected based on their diverse range of complexities. We also fine-tune BERT on our dataset to analyze the impact of fine-tuning on anti-autistic ableist speech classification.
4.2.1 Logistic Regression
We compare the performance of logistic regression (LR) supervised machine learning (ML) models to the majority of labels assigned by trained annotators. First, we use Bag of Words (BoW) as the features of the LR model. We run LR using 80%-20% train-test split of the entire dataset. These results are included in Table 5. We also compare model results with majority label downsampling since target sentences labeled with anti-autistic ableism only represent 10% of the dataset. We repeat the same LR analysis using Bag of N-Grams for both $n=2$ and $n=3$ (bi-grams and tri-grams).
Prec. Recall F1
LR 0.67 0.12 0.20
Table 5: The performance of Logistic Regression with Bag of Words.
4.2.2 BERT
We use the BERT Devlin et al. (2019) large uncased model that contains 335M parameters, again using 80% of the data for training and 20% for testing. We used a linear layer on the pooled output of the model for classification. We obtained results with BERT after fine-tuning with the AUTALIC dataset, as shown in Table 6. We conducted this experiment using the test set as a validation set to choose the best hyperparameters, as we wanted to obtain the best-case-scenario performance for a fine-tuned BERT Large on this task.
4.2.3 LLMs
We prompt GPT-4o [^5] https://openai.com/index/hello-gpt-4o/ and the instruction tuned variants of Llama-3-8B Dubey et al. (2024), Aya-23-8b Aryabumi et al. (2024), Gemma-2-2B Team et al. (2024), Mistral-7B Jiang et al. (2023), Phi-3-Mini Abdin et al. (2024) to classify the sentences in our dataset, and we adjust the prompts to compare each LLM’s performance using person-first or identity-first language. We keep the default parameters for each LLM to maintain consistency, and prompt them with the following questions:
  1. 1.
    Is this sentence ableist toward people with autism?
  2. 2.
    Is this sentence anti-autistic?
  3. 3.
    Is this sentence ableist toward autistic people?
We include each sentence from AUTALIC after the aforementioned questions in our full prompt. In addition, we provide preceding and following context for each target sentence to the LLM to mimic the level of the information supplied to human annotators.
Figure 4: The Cohen’s Kappa scores of each LLM comparing the agreement with human annotators and other LLMs after minor prompt engineering testing person-first or identity-first prompts.
4.3 Experimental Results
4.3.1 Fine-Tuning
Our experiments reveal that utilizing BERT for this classification task can lead to high rates of censorship. As BERT is not pre-trained with instructions unlike LLMs, we obtain its results after fine-tuning on AUTALIC. BERT has a high rate of false positives as nearly 2/3rds of sentences it considers ableist are not ableist according to our human annotators.
4.3.2 Human-LLM Alignment
Our assessment reveals that LLMs have low levels of alignment with human perspectives and the perspectives of other LLMs, which makes them unreliable agents for such classification tasks. We assess this alignment through a measurement using Cohen’s Kappa scores. The scores shown in Figure  4 indicate that GPT-4o had higher levels of alignment with our human-annotated dataset than any other LLM ($k=0.28$). Overall, the LLMs had low levels of agreement with human perspectives ($M=0.1029,SD=0.1096$), and even lower agreement with each other, as Phi had a worse-than-chance agreement with Gemma and Llama.
BERT Prec. Recall F1
Fine-Tuned 0.33 0.02 0.04
Table 6: The performance of BERT after fine-tuning on our dataset.
4.3.3 Safety Checks
Various LLMs have built-in safety checks, which cause them to give an invalid response, and this automatic censorship makes it difficult to employ them for the task of anti-autistic ableist speech classification. For example, Gemma has built-in safety checks that prevent it from labeling sentences that contain explicit slurs such as r*tard. Interestingly, Phi was the only LLM that did not classify certain sentences that mentioned video games without explicitly referencing autism. For example, posts discussing the Read Only Memories: NEURODIVER video game were included in our dataset as neurodiver* was one of our search keywords. These sentences simply included the titles of video games and the dates, organizations, or people affiliated with them. Notably, these safety checks only occurred when the models were prompted to detect hate speech before classifying a specific sentence. This alludes to LLMs not having an understanding of anti-autistic language as hate speech without specific instruction.
4.3.4 Understanding Ableist vs. Anti-Autistic Speech
Our results with prompt engineering reveal that anti-autistic ableism is too abstract of a concept for LLMs to recognize, providing empirical evidence that their current reasoning abilities are not inclusive of the perspectives of autistic people. LLMs struggle with identifying anti-autistic speech regardless of the terminology used, further indicating they are unreliable agents for data annotation tasks. Table 7 displays the results of prompt engineering using either person-first or identity-first language and using ’anti-autistic’ as another identity-centered way of describing this form of ableism. These results show that LLMs struggle with conceptualizing ableism directed at autistic people and its semantics. However, we can see that GPT-4o performed considerably better than the other LLMs we tested. One key difference between GPT-4o and other LLMs is the number of parameters. Based on the performance of GPT-4o on benchmarks and our understanding of scaling laws in regards to model size Kaplan et al. (2020), we estimate that GPT-4o has a far greater number of parameters than any other LLM we tested. This suggests that for larger model sizes, LLMs can begin to identify anti-autistic speech.
Model Prec. Recall F1
GPT-4o 0.25 0.61 0.36
Llama3-8B 0.12 0.88 0.21
Aya-23-8B 0.15 0.66 0.24
Gemma-2-2B 0.11 0.91 0.22
Mistral-7B 0.17 0.65 0.24
Phi-3-Mini 0.15 0.43 0.22
Table 7: The mean performance of pre-trained LLMs with prompt engineering switching from PFL to IFL.
4.4 Discussion
Our experiments demonstrate the importance of AUTALIC in aligning LLM performance to human expectations in the contexts of autism inclusion and ableist speech classification. Through our experiments, we provide empirical evidence of the current limitations of using LLMs to identify expressions of anti-autistic ableism. These limitations include: a misalignment with human perspectives, a lack of understanding of the concept of anti-autistic ableism, and censorship arising from safety checks. Each of these limitations adds to the challenge of utilizing LLMs as reliable agents for such tasks.
5 Conclusion
In this paper, we introduced AUTALIC, the first benchmark dataset focused specifically on the detection of anti-autistic ableist language in context. Through the collection and annotation of 2,400 sentences from Reddit, we aim to fill a critical gap in current NLP research, which has largely overlooked the nuanced and context-dependent nature of ableist speech targeting autistic individuals. Our dataset not only captures key contextual elements but also incorporates a comprehensive annotation process led by trained annotators with an understanding of the autistic community, ensuring higher reliability and relevance.
Our experiments demonstrate the significant limitations of existing language models, including LLMs and traditional classifiers, in effectively identifying anti-autistic ableist speech. Standard pre-trained models such as showed poor performance, reinforcing the need for specialized fine-tuning. However, even after fine-tuning BERT on the AUTALIC dataset, our experiments reveal a high rate of false positives, which can lead to unfair censorship if BERT is employed for this task. Additionally, our results reveal that even state-of-the-art LLMs struggle with implicit bias and exhibit low agreement with human annotators on this task, further emphasizing the challenges of detecting subtle forms of ableism using generic models.
Looking forward, AUTALIC paves the way for significant advancements in content moderation systems, hate speech detection models, and research on ableism and neurodiversity. We envision this dataset as a cornerstone for future work in addressing bias against autistic individuals and fostering a more inclusive digital environment. By sharing this resource with the broader research community, we aim to catalyze the development of more equitable NLP systems that better serve underrepresented and marginalized groups.
6 Limitations
Despite doing our best to include a variety of ableist language against autistic people, our dataset can still show some selection bias Ousidhoum et al. (2020) as we relied on keywords and specific social media threads to collect our data.
We acknowledge that our work centers Western perspectives in defining and understanding autism Anderson-Chavarria (2022). Therefore, we do not claim that AUTALIC is generalizable across languages and cultures as anti-autistic and ableist speech may manifest and be perceived differently.
Our dataset only contains sentences from Reddit, and subreddits focusing on autism such as r/Aspergers, r/Autism, and r/AutismInWomen were among the most popular subreddits the posts were published in. Our search criteria included words like r*tard which is a slur broadly used against many intellectual disabilities and not specifically focused on autism. As well, since we broadened our search criteria to include variations of neurodiver* such as ’neurodiversity’, ’neurodiverse’, and ’neurodivergent’, the search inevitably yielded posts discussing unrelated topics such as the NEURODIVER video game. The majority of these sentences were also published in 2023.
7 Ethics Statement
All the sentences in our dataset are publicly available, and we follow the methodologies of prior work in our data collection process  Atuhurra and Kamigaito (2024). We received IRB approval from our university’s review board and identified volunteer annotators through our association with various academic groups. Due to the sensitive nature of the content, we provided trigger warnings to our annotators and allowed them the freedom to annotate at their own pace. Additionally, we connected the members of each annotation team to enable discussions on the content and annotation process as needed.
While our dataset and citations will be made available to the academic community, commercial use of the dataset is not allowed due to the size and nature of the data. As our knowledge of anti-autistic ableism continues to evolve, AUTALIC’s classification might become outdated. While we will be adding disclaimers if needed to reflect these changes, we still encourage researchers building upon our work to stay updated on the latest semantics by referring to the perspectives of autistic scholars, activists, and organizations.

### v3

@@
 Yet, detecting anti-autistic ableist language remains underexplored, with existing NLP tools often failing to capture its nuanced expressions.
-We present AUTALIC, the first benchmark dataset dedicated to the detection of anti-autistic ableist language in context, addressing a significant gap in the field.
-The dataset comprises 2,400 autism-related sentences collected from Reddit, accompanied by surrounding context, and is annotated by trained experts with backgrounds in neurodiversity.
-Our comprehensive evaluation reveals that current language models, including state-of-the-art LLMs, struggle to reliably identify anti-autistic ableism and align with human judgments, underscoring their limitations in this domain.
-We publicly release AUTALIC along with the individual annotations which serve as a valuable resource to researchers working on ableism, neurodiversity, and also studying disagreements in annotation tasks.
+We present Autalic, the first dataset dedicated to the detection of anti-autistic ableist language in context, addressing a significant gap in the field.
+Autalic comprises 2,400 autism-related sentences collected from Reddit, accompanied by surrounding context, and annotated by trained experts with backgrounds in neurodiversity.
+Our comprehensive evaluation reveals that current language models, including state-of-the-art LLMs, struggle both to reliably identify anti-autistic ableism and to align with human judgments, underscoring their limitations in this domain.
+We publicly release Autalic along with the individual annotations.
 This dataset serves as a crucial step towards developing more inclusive and context-aware NLP systems that better reflect diverse perspectives.
@@
 Naba Rizvi1, Harper Strickland1, Daniel Gitelman1, Tristan Cooper1, Alexis Morales-Flores1, Michael Golden1, Aekta Kallepalli1, Akshat Alurkar1, Haaset Owens1, Saleha Ahmedi1, Isha Khirwadkar1, Imani Munyaka1, Nedjma Ousidhoum2 1 University of California, San Diego, 2 Cardiff University, Correspondence: nrizvi@ucsd.edu
+Trigger warning: this paper contains ableist language including explicit slurs and references to violence.
+1 Introduction
 Figure 1: The example illustrates the importance of labeling sentences in context.
@@
 Adding the surrounding sentences, as shown on the right, provides context revealing the original poster’s reference to the debunked vaccine-autism stereotype, which is tied to anti-autistic stigma Mann (2019); Davidson (2017).
-Trigger warning: this paper contains examples of anti-autistic ableist language including explicit slurs and references to violence.
-1 Introduction
-There are several critical frameworks used to define and understand disability, including autism Lawson and Beckett (2021).
-While these frameworks are ever-evolving, the medical model framework which defines disability as a “disease” is one of the most commonly applied in computer science research when discussing disability in general and autism specifically Rizvi et al.
+There are several critical frameworks used to define autism Lawson and Beckett (2021), including the medical model, which defines disability as a “disease” and is one of the most widely used in computer science research focusing on autism Rizvi et al.
 (2024); Spiel et al.
@@
 (2022).
-This framework defines autism as a deficit of skills such as the Theory of Mind, and its applications in technology research largely focus on providing diagnosis and treatment to autistic people Baron-Cohen (1997); Begum et al.
+Since this framework defines autism as a deficit of skills, its applications in technology research largely focus on providing diagnosis and treatment to autistic people Baron-Cohen (1997); Begum et al.
 (2016); Rizvi et al.
@@
 (2019b).
-However, this framework has been criticized for creating a power imbalance between medical professionals and the autistic community as it promotes the belief that a formal diagnosis is necessary to validate the autistic identity Bennett and Keyes (2020).
 This belief also posits neurotypical behaviors as the “norm” and autism as a “deficit” of these norms, thereby promoting neuronormativity instead of neurodiversity, which views all neurotypes as valid forms of human diversity Bottema-Beutel et al.
-(2021); Walker (2014).
-In contrast, prior work has shown viewing autism as an identity may increase the psychological well-being of autistic individuals and lower their social anxiety Cooper et al.
-(2023).
-Although the language used to describe autism varies, prior studies with autistic American adults found 87% prefer identity-first language over person-first language Taboas et al.
-(2023).
-Person-First Language (PFL) centers the person (i.e.
-person with autism), while Identity-First Language (IFL) centers the identity (i.e.
-autistic person) Taboas et al.
-(2023).
-Prior work examining ableist speech directed at content creators identified 11 types of “ableist hate and harassment" that they categorized as slurs and derogatory language, violent and eugenics-related speech, questioning ability and denying access, mocking and invalidation of disability identity, and objectifying the disabled body Heung et al.
-(2024).
-Additionally, researchers have identified emerging terms such as “weaponized autism" that are gaining popularity on alt-right social media platforms, further reflecting the ever-evolving nature of negative societal attitudes towards autism Welch et al.
-(2023).
+(2021a); Walker (2014).
+To improve the alignment of AI research with neurodiversity, we present Autalic, a dataset of 2,400 autism-related sentences that we collect from various communities on Reddit, along with the original context of 2,014 immediately preceding and 2,400 following sentences.
+We aim to fill a critical gap in current NLP research, which has largely overlooked the nuanced and context-dependent nature of ableist speech targeting autistic individuals.
+Our dataset not only captures key contextual elements but also incorporates a comprehensive annotation process led by trained annotators with an understanding of the autistic community, ensuring higher reliability and relevance.
+Our final dataset contains all of the labels to capture the nuances in human perspectives, and allows Autalic to serve as a resource for researchers studying anti-autistic ableist speech, neurodiversity, or disagreements in general.
+Through a series of experiments with classical models and 4 LLMs, we find empirical evidence highlighting the difficulty of this task, and that LLMs are not reliable agents for such annotations.
+Our evaluations indicate that reasoning LLMs have the most consistent scores regardless of the language used in the prompt, thereby indicating a more thorough understanding of the different ways anti-autistic speech may be identified or may manifest in text.
+We find that in-context learning examples provide mixed results in helping improve the task comprehension among LLMs.
+2 Related Work
 Anti-autistic ableist language can be diverse in scope.
 It may include perpetuating stereotypes, using offensive language and slurs, or centering non-autistic people over the perspectives of autistic people Bottema-Beutel et al.
-(2021); Rizvi et al.
+(2021a); Rizvi et al.
 (2024); Darazsdi and Bialka (2023).
@@
 (2022).
-Additionally, no datasets specifically focus on anti-autistic speech classification, and only 3 of the 23 datasets for bias evaluation in LLMs focus on disability Gallegos et al.
+Although the language used to describe autism varies, prior studies with autistic American adults found 87% prefer identity-first language over person-first language Taboas et al.
 (2023).
-Prior work has also found that LLMs lack an acknowledgment of context, which leads to higher rates of false positives when classifying ableist speech Phutane et al.
+Person-First Language (PFL) centers the person (e.g.
+“person with autism”), while Identity-First Language (IFL) centers the identity (e.g.
+“autistic person”) Taboas et al.
+(2023).
+Supporting this finding, other researchers have found that viewing autism as an identity may increase the psychological well-being of autistic individuals and lower their social anxiety Cooper et al.
+(2023).
+Ableist language online varies, may manifest in different ways and is ever-evolving  Heung et al.
+(2024); Welch et al.
+(2023).
+However, toxic language datasets focusing on hate speech and abusive language have often addressed disability in general terms but have not explicitly focused on autism ElSherief et al.
+(2018); Ousidhoum et al.
+(2019).
+To our knowledge, there are no previous datasets specifically focused on anti-autistic speech classification, and only 3 of the 23 datasets for bias evaluation in LLMs focus on disability Gallegos et al.
+(2024).
+LLMs may be limited in that they lack an acknowledgment of context, which leads to higher rates of false positives when classifying ableist speech Phutane et al.
 (2024).
@@
 (2024).
-Our work addresses these gaps in existing research by examining the performance of our dataset in the following domains for the task of anti-autistic hate speech classification:
-  1. 1.
-    The impact our dataset has on fine-tuning classifiers
-  2. 2.
-    The alignment of LLMs with the perspectives of trained human annotators
-  3. 3.
-    The extent to which in-context learning improves human-LLM alignment.
-2 Related Work
-Prior studies have also collected insights on anti-autistic language from the parents of autistic children Allouch et al.
-(2018, 2019).
-However, toxic language datasets focusing on hate speech and abusive language have often addressed disability in general terms but have not explicitly focused on autism ElSherief et al.
-(2018); Ousidhoum et al.
-(2019).
-Therefore, toxic language detection models, including LLM-based models, have been found to exhibit strong negative biases toward disabilities by classifying any disability-related text as toxic Narayanan Venkit et al.
+Toxic language detection models, including LLM-based models, have been found to exhibit strong negative biases toward disabilities by classifying any disability-related text as toxic Narayanan Venkit et al.
 (2023).
@@
 3 Autalic
-This section discusses the process of compiling and annotating the AUTALIC dataset.
-We collected relevant sentences containing autism-related keywords from Reddit using the methods described in Section 3.1.
+To build Autalic, we collected relevant sentences containing autism-related keywords from Reddit using the methods described in Section 3.1.
 The collected sentences were labeled by trained annotators, as discussed in Section 3.2.
 3.1 Data Collection
-Here, we present details from our data collection process, which include identifying our methodology for gathering representative sentences related to autism, curating the data obtained through our search parameters, and summarizing the resulting AUTALIC dataset.
+We identify a methodology for curating sentences related to autism similar to prior datasets by collecting English-language sentences from Reddit  Overbay et al.
+(2023); Wadhwa et al.
+(2023); Njoo et al.
+(2023); Antoniak et al.
+(2024); Sabri et al.
+(2024); Allein and Moens (2024); Buz et al.
+(2024).
+The limitations of this method are detailed in Section  6.
 3.1.1 Data Collection Criteria
-First, we identified potential data sources and selected Reddit as the preferred source based on its popularity, focus on text-based content, and fewer API restrictions than X at the time of our data collection.
-We searched for keywords using the default search settings on Reddit, which filters posts based on relevancy by prioritizing rare words in the search query, the age of the post, and the amount of likes and comments it has   [^1] https://support.reddithelp.com/hc/en-us/articles/19695706914196-What-filters-and-sorts-are-available.
-Table 1 displays the keywords we searched for on Reddit and the final number of sentences with each keyword in our dataset.
+We select Reddit as our source based on its popularity, focus on text-based content, and fewer API restrictions than X at the time of our data collection in January 2024.
+We search for keywords using the default search settings, which filters posts based on relevancy by prioritizing rare words in the search query, the age of the post, and the amount of likes and comments it has.
+[^1] https://support.reddithelp.com/hc/en-us/articles/19695706914196-What-filters-and-sorts-are-available The search terms include “autis*”, “ASD”, “aspergers”, and “disabilit*”; the full list is available in the Appendix, Table 4.
+Figure 2: An example of a sentence from our dataset. The search keyword is shown in red, while the word in blue is an example of a word defined in our glossary.
+We use the identified search terms to collect the target sentence instance containing our keywords, to be labeled by the annotators, and the sentences preceding or following target instances to provide additional context.
+We collect 2,400 target sentences, with 2,014 preceding and 2,400 following them.
+Finally, we split our dataset into three parts by randomly selecting and assigning 800 unique target sentences to create three segments that were each annotated by a group of three annotators.
+The average number of likes on each post included in the Autalic dataset is 1,611.59.
+Table 1 details the subreddits from which the most significant number of sentences were extracted from.
+With the exception of r/AmITheAsshole, all of the other subreddits are autism-related.
+SubReddit Sentence Count
+r/Aspergers 116
+r/Autism 88
+r/AmITheAsshole 39
+r/AutisminWomen 37
+r/AuDHDWomen 24
+Table 1: The subreddits with the most sentences included in the Autalic dataset and the number of sentences extracted from each.
+3.1.2 Data Curation
+As some of the identified keywords may appear in other contexts, we perform an exact word search for the acronyms to ensure unrelated words that might contain our acronyms are excluded from the search as they go beyond the scope of our dataset.
+For example, we searched for “applied behavioral analysis” and a case-sensitive search for “ABA”, which is a form of therapy intended to minimize autistic behaviors such as stimming (which is often used for self-soothing) Sandoval-Norton et al.
+(2019).
+Similarly, we exclude any posts that are not written in English using the Python package langdetect and posts that contain images, videos, or links.
+[^2] https://pypi.org/project/langdetect/
+3.1.3 Final Dataset
+Our final dataset includes 2,400 sentences from 192 different subreddits. To protect our annotators’ privacy, we have anonymized individual label selections.
+While nearly a quarter of the posts in our dataset were published in 2023, the range of publication years is 2013-2024.
+Figure 2 shows an example of a sentence from our dataset that uses both a search keyword and a word defined in our glossary described in Section 3.2.2.
+3.2 Data Annotation
+3.2.1 Annotator Selection
+We recruit nine upper level undergraduate researchers as volunteer annotators and randomly assign them to annotate different segments of the dataset.
+We ensure that their involvement in our annotation process is voluntary, informed, and mutually beneficial.
+In particular, some participants choose to volunteer because they care deeply about the subject matter and wish to contribute to improving AI for autistic people.
+We select participants for whom this collaboration would provide relevant and valuable professional experience, grant them opportunities to engage in other aspects of the research, and provide mentorship and authorship recognition in accordance with the ACL guidelines.
+We prioritize the well-being and autonomy of our annotators by providing full disclosure of the research process and subject material prior to their participation.
+We supply relevant trigger warnings, discuss the nature of the content in detail during an orientation session, and allow annotators to make an informed decision about whether they wish to proceed.
+We make them aware that they are free to withdraw from the study at any point.
+Our annotators are US-based, culturally diverse, and include people who grew up outside the United States.
+They are all fluent in English.
+Four of our annotators are gender minorities, and at least three self-identify as neurodivergent.
+Although we ensure the annotators were from diverse backgrounds during our recruitment process, due to the collaborative nature of our annotation process, we do not share the individual details of their identities.
+We also note that any personally identifiable information was destroyed upon the conclusion of our analysis and not shared outside of our research team.
+3.2.2 Annotator Training
+We provide a virtual orientation to all annotators explaining the history of anti-autistic ableism, examples of contemporary anti-autistic discrimination, and a brief overview of the annotation task.
+The orientation begins with a discussion of the medical model approach to autism and its link to the Nazi eugenics program Waltz (2008); Sheffer (2018).
+We define neuronormativity as the belief that the neurotypical brain is “normal” and other neurotypes are deficient in neurotypicality Wise (2023).
+We dive deeper into the medical model by discussing its impact on the self-perceptions and inclusion of autistic people in our society, such as an increase in suicidal ideation and social isolation among autistic people who mask or hide their autistic traits  Cassidy et al.
+(2014, 2018).
+Then, we cover the shifts in perspectives that emerged due to disability rights activism  Rowland (2015); Cutler (2019), and define neurodiversity as the belief that all neurotypes are valid forms of human diversity  Walker (2014).
+To explain the annotation task, we provide examples of sentences similar to what they may encounter while annotating.
+For example, we discuss how the inclusion of “at least” alters the connotations of the following sentence:
+At least I am not autistic.
+With just a minor change, the sentence can have an ableist connotation as it implies relief in knowing one is not autistic, as if it is shameful or wrong.
+We also introduce our glossary to the annotators as a dynamic resource that can be altered as needed.
+This glossary contains words that may appear in autism discourse online that may not be commonly known to others.
+These include medical acronyms, slang, and references to organizations and resources commonly affiliated with the autistic community (such as Autism Speaks).
+An excerpt of our glossary is available in our Appendix.
+We conclude our orientation by providing a brief tutorial video demonstrating how to run the script that will guide each annotator through the annotation task.
+3.2.3 Data Labeling
+After completing the training, we assign each of the three segments of the dataset to three randomly selected annotators.
+Each annotator is assigned 800 unique sentences, with a goal of completing 200 annotations each week over four weeks.
+Annotators select from three possible labels for each sentence: “Ableist,” “Not Ableist,” or “Needs More Context.”
+Ableist (1): We ask our annotators to select this label if a sentence contains ableist sentiments as defined by the Center for Disability Rights: “Ableism is a set of beliefs or practices that devalue and discriminate against people with physical, intellectual, or psychiatric disabilities and often rests on the assumption that disabled people need to be ‘fixed’ in one form or the other.” [^3] https://cdrnys.org/blog/uncategorized/ableism/
+Not Ableist (0): Annotators select this label for sentences that describe positive or neutral behaviors and attitudes regarding autism, or posts written by an autistic person reaching out for help and support.
+This includes individuals using medical terminology in a personal context (e.g., “I need therapy”), intra-community discussions, and general discussions of medical processes (unrelated to neurodivergence).
+Some examples of statements labeled as not ableist are: “I am autistic”, “As an autistic person, I think…”
+Needs More Context (-1): This label is used for sentences an annotator is unable to definitively categorize as ableist or not ableist even with the contextual sentences provided.
+This category includes text that is entirely unrelated to disabilities or remains ambiguous without additional context.
+The number of times our annotators assign each label is detailed in Table 2.
+To calculate our agreement scores, we consolidated labels -1 and 0 together based on feedback from our annotators that unrelated sentences needing more context could be classified as not anti-autistic in a purely granular classification.
+Label Definition Count
+-1 unrelated to autism or needs more context 595
+0 not ableist 5,582
+1 ableist 1,023
+Table 2: An explanation of the labels used in our classification task and the resulting counts of each label from all 9 annotators combined.
+While we use the majority label as the ground truth in our analysis, in our public dataset, we will be releasing the individual labels from each annotator due to a growing interest in embracing disagreements for such classification tasks in NLP Leonardelli et al.
+(2021); Kralj Novak et al.
+(2022); Pavlick and Kwiatkowski (2019); Plank (2022); Plank et al.
+(2014).
+Our dataset contains 2,400 sentences labeled as containing anti-autistic ableist language or not.
+The labels are obtained by calculating the mode from the three annotators of each data segment.
+Using this methodology, 242 target sentences contain examples of anti-autistic ableist language (10% of total), and 2160 sentences do not (90% of total).
+3.2.4 Providing Context
+While we provide additional sentences for context, the annotators are instructed to annotate the target sentence exclusively and only refer to the other sentences for additional context, such as determining whether the sentence is part of an intra-community discussion or the use of figurative speech (i.e.
+sarcasm).
+Figure 1 provides an example of a target sentence in context.
+In this example, it is difficult to determine whether or not the writer had ableist intent, as it can be interpreted in multiple ways.
+For example, they can be critiquing the medical model, as many autistic activists do, thereby making it non-ableist.
+Or they could be genuinely promoting ableist misrepresentations.
+The contextual sentences help the annotators better understand the writer’s intent.
+With these sentences, it is apparent that the writer is referring to the harmful and widely discredited association of vaccines with autism, which not only promotes anti-autistic ableism in society but also puts people’s lives at risk by spreading disinformation about the benefits and harms of life-saving vaccines Gabis et al.
+(2022); Taylor et al.
+(2014); Hotez (2021).
+Throughout the annotation process, annotators can edit previous annotations based on new knowledge to account for changes in language usage and connotations and the annotators’ dynamic understanding of ableism.
+3.2.5 Disagreements
+The average Fleiss’s Kappa scores are 0.25.
+This score underlines the difficulty of our classification task, which is apparent from the findings of prior works Ousidhoum et al.
+(2019), including a quantitative assessment of tag confusions that found the majority of disagreements are due to linguistically debatable cases rather than errors in annotation  Plank et al.
+(2014).
+Examples of such cases are provided in our Appendix.
+We analyze the sentences with the highest levels of disagreement in our dataset. In 100 of these posts, we observe:
+  1. 1.
+    a tendency to use the medical model terminology or stereotypes (n=48)
+  2. 2.
+    a need for additional context beyond the sentences we provided
+Figure 3: An example of a sentence from our dataset, shown in red, with a high level of disagreement among annotators.
+Figure 3 contains an example of a sentence with a high disagreement among our annotators.
+While functioning labels are considered ableist due to their eugenicist approach of categorizing autistic people based on their perceived economic value  De Hooge (2019), it is difficult to determine whether the original poster is autistic or not.
+The context is important here as classifying a sentence such as this as “ableist” can lead to unfair censorship if the original poster is a self-diagnosed autistic person seeking advice.
+Therefore, these sentences were ultimately classified as “not ableist” in Autalic.
+Our analysis reveals a moderately strong negative correlation between the task completion time and agreement with other annotators that is statistically significant ($R=-0.644$, $p$-value: $0.0096$).
+This highlights the importance of our orientation as we provided it simultaneously to the annotators.
+The annotators who completed their task immediately after our orientation had higher agreement.
+4 Experiments
+4.1 Experimental Setup
+We test the performance of different types of models on our dataset, including LLMs.
+These models are selected based on their diverse range of complexities.
+We also fine-tune BERT Devlin (2018) on our dataset to test classification.
+4.1.1 Models
+We use logistic regression (LR) with Bag of Words as the features, and using 80%-20% train-test split of the Autalic dataset.
+Fine-tuned BERT
+We also utilize fine-tuned BERT Devlin et al.
+(2019) as a baseline for LLM experiment results, using an 80%-20% train-test split.
+The F1 scores for LR and BERT (both pretrained and fine-tuned) are presented in Table 3 alongside all LLM results.
+4.1.2 Prompting LLMs
+With our baseline established, we use the LLMs Gemma2  Team et al.
+(2024), Mistral  Jiang et al.
+(2023), Llama3  Dubey et al.
+(2024), and DeepSeek  Guo et al.
+(2025) to classify the sentences in our dataset, and we adjust the prompts to compare each LLM’s performance.
+Due to limitations in computational resources, we are unable to fine-tune the LLMs.
+Prompts
+Due to the ever-evolving nature of language and variations in preferences among autistic individuals  Taboas et al.
+(2023), we use three different kinds of prompts to measure the consistency LLMs have in their understanding of anti-autistic ableist speech.
+These prompts include person-first (i.e.
+’people with autism’), identity-first (i.e.
+’autistic people’), and conceptual (i.e.
+anti-autistic) language.
+We keep the default parameters for each LLM to maintain consistency, and prompt them with the following questions:
+For each target sentence, respond to the following questions with 0 for no or 1 for yes. Refer to the preceding and following sentences if more context is needed.
+  • Is this sentence ableist toward people with autism?
+  • Is this sentence anti-autistic?
+  • Is this sentence ableist toward autistic people?
+We include each sentence from Autalic after the aforementioned questions in our full prompt.
+In addition, we provide preceding and following context for each target sentence to the LLM to mimic the level of the information supplied to human annotators.
+We run two sets of experiments with each LLM: one that uses zero-shot prompting, and another containing engineered prompts for in-context learning verbatim from the definitions and examples provided in our annotator orientation (Appendix Section A.2).
+Baselines
+Model Result PT FT
+LR 0.20
+BERT 0.43 0.90
+Simple Prompting
+LLM PFL IFL AA
+Gemma2 0.23 0.19 0.33
+Mistral 0.28 0.27 0.34
+Llama3 0.09 0.10 0.15
+DeepSeek 0.58 0.57 0.59
+In-Context Learning
+Gemma2 0.25 0.24 0.34
+Mistral 0.31 0.24 0.34
+Llama3 0.14 0.14 0.11
+DeepSeek 0.55 0.56 0.55
+Table 3: The F1 scores of various models using person-first (PFL), identify-first (IFL), and conceptual anti-autistic (AA) prompts with and without in-context learning examples for each LLM.
+The best scores for each model are in bold.
+4.2 Experimental Results
+4.2.1 Fine-Tuning
+Our experiments reveal that utilizing BERT for this classification task can lead to high rates of censorship.
+As BERT (unlike the other LLMs tested) is not pre-trained with instructions, we obtain its results after fine-tuning on Autalic.
+While the pre-trained BERT model showed poor performance indicating that it was ineffective at identifying anti-autistic ableist speech, after fine-tuning on Autalic, the model’s performance improved dramatically across all metrics, indicating that it performs better at predicting ableist speech correctly, has fewer false positives, and has a higher sensitivity to recognizing ableist speech.
+4.2.2 Human-LLM Alignment
+Our assessment reveals that LLMs have low levels of alignment with human perspectives and the perspectives of other LLMs, which makes them unreliable agents for such classification tasks.
+We assess this alignment through a measurement using Cohen’s Kappa scores.
+The scores shown in Figure 4 indicate the highest level of alignment was demonstrated between Gemma2 and Mistral ($k=0.34$).
+No LLM demonstrated alignment with our human-annotated dataset, although DeepSeek’s alignment was notably higher than the others.
+Overall, the LLMs had low levels of agreement with human perspectives ($M=0.091,SD=0.110$).
+This indicates that LLMs with less than 10 billion parameters struggle with the task of classifying anti-autistic ableist language, even when provided with in-context examples.
+4.2.3 In-Context Learning
+After providing the in-context learning examples, Llama3 (+22.96%) and Gemma2 (+12.68%) display the biggest relative improvement in F-1 scores, indicating that both models benefit from the examples.
+In particular, we find our ICL examples resulted in large relative improvements in consistency of scores regardless of the language used in the prompts for each LLM.
+For example, providing Llama with examples helped decrease its relative change in F1 scores from 67.49% to 17.4% when switching from PFL to conceptual prompts, indicating a better understanding of the connection between anti-autistic ableism and ableism toward autistic people.
+Figure 4: The mean Cohen’s Kappa scores of each LLM comparing the agreement with human annotators and other LLMs.
+4.2.4 Understanding Ableist and Anti-Autistic Speech
+Our results with prompt engineering reveal that anti-autistic ableism is too abstract of a concept for LLMs to recognize, providing empirical evidence that their current reasoning abilities are not inclusive of the perspectives of autistic people.
+LLMs struggle with identifying anti-autistic speech regardless of the terminology used, further indicating they are unreliable agents for data annotation tasks.
+Table 3 displays the results of prompt engineering using either person-first or identity-first language and using “anti-autistic” to describe this form of ableism in more conceptual terms.
+Notably, switching from PFL or IFL to conceptual language in the prompts resulted in the largest relative changes in scores.
+For instance, prompting Llama with "is this sentence anti-autistic?" instead of "is this sentence ableist toward people with autism?" resulted in a relative increase of 67.49%.
+These results show that LLMs struggle with understanding that “anti-autistic ableism" and “ableism toward autistic people" refer to the same phenomenon.
+Even after providing the ICL examples, the relative change in F1 scores between different prompts was as high as 32.66% for Gemma2.
+Interestingly, DeepSeek, the only reasoning LLM we test in our study, has the best results and highest consistency out of all the other LLMs.
+Although its agreement with human annotators is low ($k:0.11$), it is still double that of all other LLMs, as shown in Figure  4.
+This highlights the difficulty of this task, as more advanced reasoning is required to understand the nuances of anti-autistic ableism, including the terminology we use to describe such speech.
+4.3 Discussion
+Classifying anti-autistic ableist speech is challenging even within a Western, English-speaking context, since perceptions of what constitutes anti-autistic content differ significantly even among autistic individuals Keating et al.
+(2023).
+Factors such as personal experiences, cultural norms, and evolving discourse around autism advocacy can influence each individual’s perception of or sensitivity toward recognizing toxic speech, making it difficult to establish a consistent classification scheme Ousidhoum et al.
+(2019); Bottema-Beutel et al.
+(2021b); Taboas et al.
+(2023); Kapp et al.
+(2013).
+While developing Autalic, we standardize our definition of anti-autistic ableist speech by providing our annotators with an orientation and a glossary.
+These resources are developed in alignment with the perspectives of autistic people  Bottema-Beutel et al.
+(2021a); Taboas et al.
+(2023); Kapp et al.
+(2013).
+Our experiments demonstrate the importance of Autalic in aligning LLM performance to human expectations in the contexts of autism inclusion and ableist speech classification.
+Through our experiments, we provide empirical evidence of the current limitations of using LLMs and traditional classifiers to identify expressions of anti-autistic ableism.
+These limitations include: a misalignment with human perspectives, a lack of understanding of the concept of anti-autistic ableism, and a lack of agreement with each other even with in-context learning examples.
+Each of these limitations adds to the challenge of utilizing LLMs as reliable agents for such tasks.
+Standard pre-trained models such as showed poor performance, reinforcing the need for specialized fine-tuning.
+However, even after fine-tuning BERT on the Autalic dataset, our experiments reveal a high rate of false positives, which can lead to unfair censorship if BERT is employed for this task.
+Additionally, our results reveal that even state-of-the-art LLMs exhibit low agreement with human annotators on this task, further emphasizing the challenges of detecting subtle forms of ableism using generic models.
+DeepSeek has the best performance out of all the LLMs in our study, further demonstrating the difficulty of this task, as it is the only a reasoning-focused LLM in our study.
+5 Conclusion
+In this paper, we introduced Autalic, the first benchmark dataset focused specifically on the detection of anti-autistic ableist language in context.
+Through the collection and annotation of 2,400 sentences from Reddit, we aim to fill a critical gap in current NLP research and improve its alignment with neurodiversity.
+Looking forward, Autalic paves the way for significant advancements in content moderation systems, hate speech detection models, and research on ableism and neurodiversity.
+We envision this dataset as a cornerstone for future work in addressing bias against autistic individuals and fostering a more inclusive digital environment.
+By sharing this resource with the broader research community, we aim to catalyze the development of more equitable NLP systems that better serve underrepresented and marginalized groups.
+6 Limitations
+Despite doing our best to include a variety of ableist language against autistic people, our dataset can still show some selection bias Ousidhoum et al.
+(2020) as we relied on keywords and specific social media threads to collect our data.
+We recognize that, like other datasets presented at ACL, our work primarily reflects Western perspectives  Anderson-Chavarria (2022); Kirk et al.
+(2023); Thapa et al.
+(2022); Aoyama et al.
+(2023); Takeshita et al.
+(2024); Park and Park (2023).
+Therefore, we do not claim that Autalic is generalizable across languages and cultures as anti-autistic and ableist speech may manifest and be perceived differently.
+However, we emphasize that Autalic is the first dataset to center the perspectives of autistic people in ableist speech detection, and move away from the tendency of AI to misclassify any text related to disability as ‘toxic’  Narayanan Venkit et al.
+(2023); Heung et al.
+(2024); Van Dorpe et al.
+(2023); Hutchinson et al.
+(2020).
+Therefore, it can serve as a reference for future work exploring ableist speech in diverse contexts such as other cultures and languages.
+Our dataset contains sentences from Reddit, and subreddits focusing on autism such as r/Aspergers, r/Autism, and r/AutismInWomen were among the most popular subreddits the posts were published in.
+Our search criteria included words like “r*tard” which is a slur broadly used against many intellectual disabilities and not specifically focused on autism.
+While we did our best to filter out irrelevant posts, since we broadened our search criteria to include variations of “neurodiver*” such as “neurodiversity”, “neurodiverse”, and “neurodivergent”, the search inevitably yielded posts discussing unrelated topics such as the NEURODIVER video game.
+The majority of these sentences were also published in 2023.
+While our dataset is small, the quality of our labels is high due to our standardized annotation protocols and resources that were developed using rigorous and iterative testing.
+7 Ethics Statement
+The data collected for our small-scale and non-commerical research is in compliance with Reddit’s API limits and policies  Reddit (2023a, b).
+All the sentences in our dataset are publicly available, and we follow the methodologies of prior work in our data collection process  Atuhurra and Kamigaito (2024).
+We specifically recruit annotators for whom this collaboration and resulting paper authorship would be mutually beneficial, and provide a comprehensive overview of the task to ensure they make an informed decision to participate.
+Some of our annotators chose to volunteer as they care deeply about neurodiversity and autism inclusion.
+We received IRB approval from our university’s review board and identified volunteer annotators through our association with various academic groups.
+Given the sensitivity of the content, we provided annotators with appropriate trigger warnings, ensuring they could work at their own pace or withdraw from the study if necessary.
+Additionally, we connected the members of each annotation team to enable discussions on the content and annotation process as needed.
+While our dataset and citations will be made available to the academic community, commercial use of the dataset is not allowed due to the size and nature of the data.
+As our knowledge of anti-autistic ableism continues to evolve, Autalic’s classification might become outdated.
+While we will be adding disclaimers if needed to reflect these changes, we still encourage researchers building upon our work to stay updated on the latest semantics by referring to the perspectives of autistic scholars, activists, and organizations.
+Refer to our Appendix for our Guidelines for Responsible Use.
+Appendix A Appendix
+A.1 Guidelines for Responsible Use
+A.1.1 Purpose and Scope
+This dataset has been curated to aid in the classification and study of anti-autistic ableist language in a U.S.
+context using text from Reddit.
+It aims to support research and educational endeavors focused on understanding, identifying, and mitigating ableist speech directed at autistic individuals, while moving away from mis-classifying any speech related to autism or disability as toxic.
+A.1.2 Applicability and Cultural Context
+U.S.-Specific Results: The language examples and classification models in this dataset are primarily reflective of usage and cultural nuances in the United States.
+As a result, the dataset and any models developed from it may not be fully accurate or generalizable for other countries or cultural contexts.
+Data Curation: The data included in this has been taken solely from posts and comments on Reddit and may not represent autism discourse on other platforms and in other contexts.
+Contact for Latest Version: Language evolves over time. For the most up-to-date version of the dataset, or on more information on when the dataset was last updated, please contact the first author.
+A.1.3 Access and Security
+Password Protection: The dataset is password-protected to prevent unauthorized or automated scraping (e.g., by bots).
+While the password is publicly available as of this publication, it may require prior approval in the future as needed to ensure reponsible use.
+Secure Storage: Users are expected to maintain secure protocols (e.g., encryption, controlled access) to prevent unauthorized sharing or leaks of the dataset.
+The dataset may not be shared without consent of the authors.
+A.1.4 Permitted Uses
+Free Use for Scientific Research: The dataset is publicly available without charge for legitimate scientific, academic, or educational research purposes, subject to the restrictions outlined below.
+Academic and Non-Profit Contexts: Users in academic, research, or non-profit institutions may incorporate the dataset into studies, presentations, or scholarly articles, provided they follow these guidelines and appropriately cite the dataset and its authors.
+A.1.5 Prohibited or Restricted Uses Commercial Use:
+Commercial use is not authorized without explicit written permission from the dataset authors.
+If you wish to incorporate the dataset into commercial products or services, you must obtain approval in advance.
+Automated Content Moderation: Using the dataset to develop or deploy automated content moderation tools is not authorized without prior approval from the authors.
+This restriction helps ensure that any moderation system is deployed ethically and with proper considerations for context and language evolution.
+A.1.6 Ethical Considerations and Privacy
+Respect for Individuals and Communities: Users must handle the dataset with an understanding of the impacts of ableist language on autistic communities.
+The dataset’s examples are provided solely for research and analysis and must not be used to perpetuate or normalize ableist attitudes, or to scrutinize or attack any individual annotators or original posters.
+This work is not intended as an ethical judgment or targeting of individuals, but rather an effort to improve AI alignment with the perspectives of autistic people.
+Citations and Acknowledgements: When publishing findings, users should cite this dataset, acknowledging the work of its authors and the communities that provided the materials or data.
+Compliance with Regulations: Researchers must comply with all relevant local, national, and international regulations and guidelines relating to data privacy and human subjects research where applicable.
+A.1.7 How to Request Approval
+Commercial or Moderation Use: If you intend to use the dataset for commercial purposes or automated content moderation, please submit a formal request, detailing:
+  • Project objectives
+  • Potential for data use and distribution
+  • Mechanisms to ensure ethical application and protection of the data
+  • The impact of the project, and its target end-users
+Contact the First Author: All requests and inquiries should be directed to the first author, as listed in the dataset documentation or project website, available here: [REDACTED FOR PRIVACY]
+A.1.8 Liability and Disclaimer
+The dataset is provided “as is,” without any guarantees regarding completeness, accuracy, or fitness for a particular purpose, especially outside of the U.S.
+context, for multi-media posts, or discussions on other platforms outside of Reddit.
+User Responsibility: Users bear the responsibility for ensuring their use complies with these guidelines, as well as any applicable laws and ethical standards.
+By accessing and using this dataset, you acknowledge that you have read and agreed to these Guidelines for Responsible Use, and that you understand the conditions under which the dataset may be utilized for your research or projects.
+A.2 Annotator Orientation
+A.2.1 Introduction
+This subsection provides an overview of the annotation orientation session conducted for Autalic.
+The goal is to ensure annotators understand the history and contemporary examples of anti-autistic ableism, the importance of neurodiversity, and the different ways in which anti-autistic speech may manifest in text.
+Given the sensitive nature of this work, annotators are advised that they may encounter discussions involving ableist language, violence, self-harm, and suicide mentions.
+A.2.2 Understanding Anti-Autistic Ableism
+Anti-autistic ableism is the discrimination and devaluation of autistic individuals based on neuronormative standards.
+A striking example includes cases where caretakers harm autistic individuals due to societal stigma  of Mourning (2024).
+The historical roots of such bias date back to Nazi-era eugenics research, where Hans Asperger categorized autistic individuals as either “useful” or “unfit,” reinforcing a harmful hierarchical perception of autism Furfaro (2018).
+Neuronormativity
+Neuronormativity is the societal belief that neurotypical cognition is the default and that neurodivergence is an abnormality  Huijg (2020).
+This belief system marginalizes autistic individuals and contributes to discrimination in various aspects of life, including education, employment, and social interactions.
+Deficit-Based Approaches and Their Harms
+Traditional medical models frame autism as a disorder requiring intervention or treatment  Kapp et al. (2013); Kapp (2019). This perspective has led to:
+  • Increased exposure to violence and self-harm risk
+  • Social exclusion and stigmatization
+  • Internalized ableism and lower self-esteem
+Benevolent Ableism
+Benevolent ableism refers to actions or attitudes that, while seemingly supportive, reinforce autistic individuals as “less than” neurotypicals  Nario-Redmond et al.
+(2019).
+Examples include organizations like Autism Speaks, which promote awareness campaigns that fail to center autistic voices  Rosenblatt (2022).
+The use of symbols such as the puzzle piece is an example of this issue, as it implies that autism is a mystery to be solved rather than a valid identity.
+A.2.3 The Neurodiversity Movement
+The neurodiversity paradigm challenges the medical model by recognizing neurological variations as a natural and valid part of human diversity  Walker (2014).
+Symbols such as the rainbow infinity sign inspired by the LGBTQ Pride flag have emerged from within the community to counter external narratives that frame autism as a deficit  Kattari et al.
+(2023).
+Community Perspectives
+Autistic individuals often reclaim language and challenge neuronormative narratives. Important considerations for annotation include:
+  • Identity-first language (e.g., “autistic person” instead of “person with autism”) is preferred by the majority of autistic adults in the United States  Taboas et al. (2023)
+  • Community-adopted terminology such as Aspie (a self-identifier used by some autistic individuals)
+A.2.4 Annotation Tasks and Procedures
+In this section, we provide an overview of the annotation task along with video examples of the process.
+Common Annotation Challenges
+Annotators should exercise careful judgment when evaluating phrases. For example:
+  • Statements such as “That’s so autistic” require contextual interpretation.
+  • The phrase “This vaccine causes autism” is categorized as ableist due to its history in promoting autism stigma.
+  • The subtle difference between “I am not autistic” and “At least I am not autistic” changes the meaning and must be carefully assessed.
+A.2.5 Ethical Considerations and AI Bias
+Challenges in Hate Speech Detection
+Research indicates that many existing AI models misclassify disability-related discourse as toxic, even when the content is neutral or positive  Narayanan Venkit et al.
+(2023); Venkit et al.
+(2022); Gadiraju et al.
+(2023a); Gama (2024).
+Specific issues include:
+  • AI models exhibit over-sensitivity to disability-related discussions, frequently labeling them as harmful.
+  • AI models are more confident in detecting ableism when using person-first language (e.g., “ableist toward autistic people”) than identity-first language (e.g., “anti-autistic”). *
+*these are results from our preliminary study
+Project Overview
+This project seeks to mitigate biases in AI hate speech detection by:
+  • Training models using annotations informed by the neurodivergent community.
+  • Ensuring that AI does not misclassify community discourse as hate speech.
+  • Recognizing the distinction between hate speech and reclaimed terminology within the autistic community.
+A.2.6 Resources
+In this section, we provide resources such as our guidelines that contain a glossary to refer to or modify as needed.
+The terms in the glossary are those commonly used in neurodiversity discourse online.
+A.2.7 Conclusion
+The annotation orientation session is designed to equip annotators with the necessary knowledge to responsibly and accurately classify anti-autistic hate speech.
+By following the annotation guidelines and considering the broader socio-historical context, annotators contribute to the development of AI models that better serve neurodivergent individuals.
+A.3 Search Keywords
+This list of terms in Table 4 were used to identify target sentences on Reddit. The number of target sentences containing each term is included.
 Word Sentence Count
@@
 Table 4: The keywords included in our Reddit search and the number of sentences containing each term in the Autalic dataset.
-We use the identified search terms detailed in Table  1 to collect two types of sentences:
-  • the target instance containing our keywords, to be labeled by the annotators,
-  • the sentences preceding or following target instances to provide additional context.
-We collected 2,500 target sentences, with 2,014 preceding and 2,500 following them.
-Finally, we split our dataset into three parts by randomly selecting and assigning 800 unique target sentences to create three segments that were each annotated by a group of three annotators.
-3.1.2 Data Curation
-As some of the identified keywords may appear in other contexts, we perform an exact word search for the acronyms to ensure unrelated words that might contain our acronyms are excluded from the search as they go beyond the scope of our dataset.
-For example, we searched for “applied behavioral analysis” and a case-sensitive search for ABA, which is a form of therapy intended to minimize autistic behaviors such as stimming (which is often used for self-soothing) Sandoval-Norton et al.
-(2019).
-Similarly, we exclude any posts that are not written in English using the Python package langdetect  [^2] https://pypi.org/project/langdetect/ and posts that contain images, videos, or links.
-3.1.3 Final Dataset
-Our final dataset includes 2,400 sentences from 192 different subreddits. To protect our annotators’ privacy, we have anonymized all labels.
-While nearly a quarter of the posts in our dataset were published in 2023, the range of publication years is 2013-2024.
-Figure 2 shows an example of a sentence from our dataset that uses both a search keyword and a word defined in our glossary described in Section 3.2.2.
-Figure 2: An example of a sentence from our dataset. The search keyword is shown in red, while the word in blue is an example of a word defined in our glossary.
-The average number of likes on each post included in the AUTALIC dataset is 1,611.59, and only 135 posts have 0-5 likes.
-Table 2 details the subreddits from which we extracted the most significant number of sentences for our dataset.
-SubReddit Sentence Count
-r/Aspergers 116
-r/Autism 88
-r/AmITheAsshole 39
-r/AutisminWomen 37
-r/AuDHDWomen 24
-Table 2: The subreddits with the most sentences included in the AUTALIC dataset and the number of sentences extracted from each.
-3.2 Data Annotation
-After obtaining sample sentences with contextual surrounding sentences, we proceeded to annotation: annotator selection, training, and data annotation.
-3.2.1 Annotator Selection
-We recruited nine annotators and randomly assigned them to annotate different segments of the dataset.
-Our annotators are undergraduate students who volunteered to annotate the data without financial compensation and were offered authorship following the ACL guidelines on authorship.
-They are culturally diverse and include people who grew up outside the United States.
-However, they are all fluent in English.
-Four of our annotators are gender minorities, and at least three self-identify as neurodivergent.
-Although we ensured the annotators were from diverse backgrounds during our recruitment process, due to the collaborative nature of our annotation process, we cannot share the individual details of their identities.
-We also note that any personally identifiable information was destroyed upon the conclusion of our analysis and not shared outside of our research team.
-3.2.2 Annotator Training
-We provided a virtual orientation to all annotators explaining the history of anti-autistic ableism, examples of contemporary anti-autistic discrimination, and a brief overview of the annotation task.
-The orientation begins with a discussion of the medical model approach to autism and its link to the Nazi eugenics program Waltz (2008); Sheffer (2018).
-We define neuronormativity as the belief that the neurotypical brain is “normal” and other neurotypes are deficient in neurotypicality Wise (2023).
-We dive deeper into the medical model by discussing its impact on the self-perceptions and inclusion of autistic people in our society, such as an increase in suicidal ideation and social isolation among autistic people who mask or hide their autistic traits  Cassidy et al.
-(2014, 2018).
-Then, we cover the shifts in perspectives that emerged due to disability rights activism  Rowland (2015); Cutler (2019), and define neurodiversity as the belief that all neurotypes are valid forms of human diversity  Walker (2014).
-To explain the annotation task, we provide examples of sentences similar to what they may encounter while annotating.
-For example, we discuss how the inclusion of "at least" alters the connotations of the follow sentence:
-At least I am not autistic.
-With just a minor change, the sentence can have an ableist connotation as it implies relief in knowing one is not autistic, as if it is shameful or wrong.
-We also introduce our glossary to the annotators as a dynamic resource that can be altered as needed.
-This glossary contains words that may appear in autism discourse online that may not be commonly known to others.
-These include medical acronyms, slang, and references to organizations and resources commonly affiliated with the autistic community (such as Autism Speaks).
-Selected examples from our glossary are shown in Table  3.
-The final version of our glossary is available can be found in this footnote   [^3] https://figshare.com/s/3850e63295ccf6fceebb .
+A.4 Glossary (Excerpt)
+Below is an example of terms in our glossary.
+To view the full glossary, contact the first-author of this paper.
+This resource was created in collaboration with our annotators to define unfamiliar terms such as slang, medical abbreviations, laws, and other specialized language that they may encounter during the annotation process.
+In total, it contains 34 words.
 Term Definition
-AuDHD A combination of ADHD and autism
-ND Neurodivergent or neurodiverse
-NT Neurotypical
-Table 3: Selected examples from our glossary detailing commonly used acronyms, abbreviations, and slang in autism discourse online. For the full glossary, see footnote 3.
-We conclude our orientation by providing a brief tutorial video demonstrating how to run the script that will guide each annotator through the annotation task.
-3.2.3 Data Labeling
-After completing the training, we assign each of the three segments of the dataset to three randomly selected annotators.
-Each annotator is assigned 800 unique sentences, with a goal of completing 200 annotations each week over four weeks.
-Annotators select from three possible labels for each sentence: “Ableist,” “Not Ableist,” or “Needs More Context.” According to the Center for Disability Rights, “Ableism is a set of beliefs or practices that devalue and discriminate against people with physical, intellectual, or psychiatric disabilities and often rests on the assumption that disabled people need to be ‘fixed’ in one form or the other.” [^4] https://cdrnys.org/blog/uncategorized/ableism/ We supplied this definition to annotators and instructed them to select the “ableist” label for any sentences that contain examples of these beliefs and practices directed towards autistic people.
-Annotators are instructed to select the “not ableist” label for sentences that describe positive behaviors and attitudes regarding autism or disabilities or may be written by an autistic person reaching out for help and support.
-This includes individuals using medical terminology in a personal context (e.g., ‘I need therapy’), discussion and suggestions from neurodivergent people (community-generated discussions), and general discussions of the medical processes (unrelated to neurodivergence).
-Some examples of statements labeled as not ableist are: “I am autistic”, “As an autistic person, I think…”
-The “Needs More Context” label is used for sentences an annotator is unable to definitively categorize as ableist or not ableist with only the surrounding sentences as context.
-This category includes text that is entirely unrelated to disabilities or remains ambiguous without additional context.
-The labels used in our annotation task are -1 for sentences that are unrelated/need more context, 0 for sentences that are not anti-autistic, and 1 for sentences that are anti-autistic.
-The number of times our annotators assigned each label in our dataset is detailed in Table 4.
-To calculate our agreement scores, we consolidated labels -1 and 0 together based on feedback from our annotators that unrelated sentences needing more context could be classified as not anti-autistic in a purely granular classification.
-Label Definition Count
--1 unrelated to autism or needs more context 595
-0 not ableist 5,582
-1 ableist 1,023
-Table 4: An explanation of the labels used in our classification task and the resulting counts of each label from all annotators combined.
-While we use the majority label as the ground truth in our analysis, in our public dataset, we will be releasing the individual labels from each annotator due to a growing interest in embracing disagreements for such classification tasks in NLP Leonardelli et al.
-(2021); Kralj Novak et al.
-(2022); Pavlick and Kwiatkowski (2019).
-3.2.4 Providing Context
-While we provide additional sentences for context, the annotators are instructed to annotate the target sentence exclusively and only refer to the other sentences for additional context, such as determining whether the sentence is part of an intra-community discussion or the use of figurative speech (i.e.
-sarcasm).
-Figure 1 provides an example of a target sentence in context.
-In this example, it is difficult to determine whether or not the writer had ableist intent, as it can be interpreted in multiple ways.
-For example, they can be critiquing the medical model, as many autistic activists do, thereby making it non-ableist.
-Or they could be genuinely promoting ableist misrepresentations.
-The contextual sentences help the annotators better understand the writer’s intent.
-With these sentences, it is apparent that the writer is referring to the harmful and widely discredited association of vaccines with autism, which not only promotes anti-autistic ableism in society but also puts people’s lives at risk by spreading disinformation about the benefits and harms of life-saving vaccines Gabis et al.
-(2022); Taylor et al.
-(2014); Hotez (2021).
-Throughout the annotation process, annotators can edit previous annotations based on new knowledge to account for changes in language usage and connotations and the annotators’ dynamic understanding of ableism.
-3.2.5 Disagreements
-The Fleiss’s Kappa scores were $0.125$ for Team 1, $0.307$ for Team 2, and $0.191$ for team 3.
-These scores underline the difficulty of our classification task, which is apparent from the findings of prior work Ousidhoum et al.
-(2019).
-We analyzed the sentences with the highest levels of disagreement in our dataset.
-We found that such posts:
-  1. 1.
-    tend to use medical model terminology or stereotypes
-  2. 2.
-    need additional context beyond the sentences we provided.
-Our dataset contains 59 of these sentences.
-Figure 3: An example of a sentence from our dataset, shown in red, with a high level of disagreement among annotators.
-Figure 3 contains an example of a sentence with a high disagreement among our annotators.
-While functioning labels are considered ableist due to their eugenicist approach of categorizing autistic people based on their perceived economic value  De Hooge (2019), it is difficult to determine whether the original poster is autistic or not.
-The context is important here as classifying a sentence such as this as ’ableist’ can lead to unfair censorship if the original poster is a self-diagnosed autistic person seeking advice.
-Therefore, these sentences were ultimately classified as ’not ableist’ in AUTALIC.
-Our analysis reveals a moderately strong negative correlation between the task completion time and agreement with other annotators that is statistically significant ($R=-0.644$, $p$-value: $0.0096$).
-This highlights the importance of our orientation as we provided it simultaneously to the annotators, and the annotators who completed their task immediately after our orientation had higher agreement.
-4 Experiments
-4.1 Data
-Our dataset contains 2,400 sentences labeled as containing anti-autistic ableist language or not.
-The labels are obtained by calculating the mode from the three annotators of each data segment.
-Using this methodology, 242 target sentences contain examples of anti-autistic ableist language (10% of total), and 2160 sentences do not (90% of total).
-4.2 Experimental Setup
-We test the performance of different types of models on our dataset, including logistic regression, BERT, and LLMs ranging from PHI 3 Mini to GPT 4.
-These models are selected based on their diverse range of complexities.
-We also fine-tune BERT on our dataset to analyze the impact of fine-tuning on anti-autistic ableist speech classification.
-4.2.1 Logistic Regression
-We compare the performance of logistic regression (LR) supervised machine learning (ML) models to the majority of labels assigned by trained annotators.
-First, we use Bag of Words (BoW) as the features of the LR model.
-We run LR using 80%-20% train-test split of the entire dataset.
-These results are included in Table 5.
-We also compare model results with majority label downsampling since target sentences labeled with anti-autistic ableism only represent 10% of the dataset.
-We repeat the same LR analysis using Bag of N-Grams for both $n=2$ and $n=3$ (bi-grams and tri-grams).
-Prec. Recall F1
-LR 0.67 0.12 0.20
-Table 5: The performance of Logistic Regression with Bag of Words.
-4.2.2 BERT
-We use the BERT Devlin et al.
-(2019) large uncased model that contains 335M parameters, again using 80% of the data for training and 20% for testing.
-We used a linear layer on the pooled output of the model for classification.
-We obtained results with BERT after fine-tuning with the AUTALIC dataset, as shown in Table 6.
-We conducted this experiment using the test set as a validation set to choose the best hyperparameters, as we wanted to obtain the best-case-scenario performance for a fine-tuned BERT Large on this task.
-4.2.3 LLMs
-We prompt GPT-4o [^5] https://openai.com/index/hello-gpt-4o/ and the instruction tuned variants of Llama-3-8B Dubey et al.
-(2024), Aya-23-8b Aryabumi et al.
-(2024), Gemma-2-2B Team et al.
-(2024), Mistral-7B Jiang et al.
-(2023), Phi-3-Mini Abdin et al.
-(2024) to classify the sentences in our dataset, and we adjust the prompts to compare each LLM’s performance using person-first or identity-first language.
-We keep the default parameters for each LLM to maintain consistency, and prompt them with the following questions:
-  1. 1.
-    Is this sentence ableist toward people with autism?
-  2. 2.
-    Is this sentence anti-autistic?
-  3. 3.
-    Is this sentence ableist toward autistic people?
-We include each sentence from AUTALIC after the aforementioned questions in our full prompt.
-In addition, we provide preceding and following context for each target sentence to the LLM to mimic the level of the information supplied to human annotators.
-Figure 4: The Cohen’s Kappa scores of each LLM comparing the agreement with human annotators and other LLMs after minor prompt engineering testing person-first or identity-first prompts.
-4.3 Experimental Results
-4.3.1 Fine-Tuning
-Our experiments reveal that utilizing BERT for this classification task can lead to high rates of censorship.
-As BERT is not pre-trained with instructions unlike LLMs, we obtain its results after fine-tuning on AUTALIC.
-BERT has a high rate of false positives as nearly 2/3rds of sentences it considers ableist are not ableist according to our human annotators.
-4.3.2 Human-LLM Alignment
-Our assessment reveals that LLMs have low levels of alignment with human perspectives and the perspectives of other LLMs, which makes them unreliable agents for such classification tasks.
-We assess this alignment through a measurement using Cohen’s Kappa scores.
-The scores shown in Figure  4 indicate that GPT-4o had higher levels of alignment with our human-annotated dataset than any other LLM ($k=0.28$).
-Overall, the LLMs had low levels of agreement with human perspectives ($M=0.1029,SD=0.1096$), and even lower agreement with each other, as Phi had a worse-than-chance agreement with Gemma and Llama.
-BERT Prec. Recall F1
-Fine-Tuned 0.33 0.02 0.04
-Table 6: The performance of BERT after fine-tuning on our dataset.
-4.3.3 Safety Checks
-Various LLMs have built-in safety checks, which cause them to give an invalid response, and this automatic censorship makes it difficult to employ them for the task of anti-autistic ableist speech classification.
-For example, Gemma has built-in safety checks that prevent it from labeling sentences that contain explicit slurs such as r*tard.
-Interestingly, Phi was the only LLM that did not classify certain sentences that mentioned video games without explicitly referencing autism.
-For example, posts discussing the Read Only Memories: NEURODIVER video game were included in our dataset as neurodiver* was one of our search keywords.
-These sentences simply included the titles of video games and the dates, organizations, or people affiliated with them.
-Notably, these safety checks only occurred when the models were prompted to detect hate speech before classifying a specific sentence.
-This alludes to LLMs not having an understanding of anti-autistic language as hate speech without specific instruction.
-4.3.4 Understanding Ableist vs. Anti-Autistic Speech
-Our results with prompt engineering reveal that anti-autistic ableism is too abstract of a concept for LLMs to recognize, providing empirical evidence that their current reasoning abilities are not inclusive of the perspectives of autistic people.
-LLMs struggle with identifying anti-autistic speech regardless of the terminology used, further indicating they are unreliable agents for data annotation tasks.
-Table 7 displays the results of prompt engineering using either person-first or identity-first language and using ’anti-autistic’ as another identity-centered way of describing this form of ableism.
-These results show that LLMs struggle with conceptualizing ableism directed at autistic people and its semantics.
-However, we can see that GPT-4o performed considerably better than the other LLMs we tested.
-One key difference between GPT-4o and other LLMs is the number of parameters.
-Based on the performance of GPT-4o on benchmarks and our understanding of scaling laws in regards to model size Kaplan et al.
-(2020), we estimate that GPT-4o has a far greater number of parameters than any other LLM we tested.
-This suggests that for larger model sizes, LLMs can begin to identify anti-autistic speech.
-Model Prec. Recall F1
-GPT-4o 0.25 0.61 0.36
-Llama3-8B 0.12 0.88 0.21
-Aya-23-8B 0.15 0.66 0.24
-Gemma-2-2B 0.11 0.91 0.22
-Mistral-7B 0.17 0.65 0.24
-Phi-3-Mini 0.15 0.43 0.22
-Table 7: The mean performance of pre-trained LLMs with prompt engineering switching from PFL to IFL.
-4.4 Discussion
-Our experiments demonstrate the importance of AUTALIC in aligning LLM performance to human expectations in the contexts of autism inclusion and ableist speech classification.
-Through our experiments, we provide empirical evidence of the current limitations of using LLMs to identify expressions of anti-autistic ableism.
-These limitations include: a misalignment with human perspectives, a lack of understanding of the concept of anti-autistic ableism, and censorship arising from safety checks.
-Each of these limitations adds to the challenge of utilizing LLMs as reliable agents for such tasks.
-5 Conclusion
-In this paper, we introduced AUTALIC, the first benchmark dataset focused specifically on the detection of anti-autistic ableist language in context.
-Through the collection and annotation of 2,400 sentences from Reddit, we aim to fill a critical gap in current NLP research, which has largely overlooked the nuanced and context-dependent nature of ableist speech targeting autistic individuals.
-Our dataset not only captures key contextual elements but also incorporates a comprehensive annotation process led by trained annotators with an understanding of the autistic community, ensuring higher reliability and relevance.
-Our experiments demonstrate the significant limitations of existing language models, including LLMs and traditional classifiers, in effectively identifying anti-autistic ableist speech.
-Standard pre-trained models such as showed poor performance, reinforcing the need for specialized fine-tuning.
-However, even after fine-tuning BERT on the AUTALIC dataset, our experiments reveal a high rate of false positives, which can lead to unfair censorship if BERT is employed for this task.
-Additionally, our results reveal that even state-of-the-art LLMs struggle with implicit bias and exhibit low agreement with human annotators on this task, further emphasizing the challenges of detecting subtle forms of ableism using generic models.
-Looking forward, AUTALIC paves the way for significant advancements in content moderation systems, hate speech detection models, and research on ableism and neurodiversity.
-We envision this dataset as a cornerstone for future work in addressing bias against autistic individuals and fostering a more inclusive digital environment.
-By sharing this resource with the broader research community, we aim to catalyze the development of more equitable NLP systems that better serve underrepresented and marginalized groups.
-6 Limitations
-Despite doing our best to include a variety of ableist language against autistic people, our dataset can still show some selection bias Ousidhoum et al.
-(2020) as we relied on keywords and specific social media threads to collect our data.
-We acknowledge that our work centers Western perspectives in defining and understanding autism Anderson-Chavarria (2022).
-Therefore, we do not claim that AUTALIC is generalizable across languages and cultures as anti-autistic and ableist speech may manifest and be perceived differently.
-Our dataset only contains sentences from Reddit, and subreddits focusing on autism such as r/Aspergers, r/Autism, and r/AutismInWomen were among the most popular subreddits the posts were published in.
-Our search criteria included words like r*tard which is a slur broadly used against many intellectual disabilities and not specifically focused on autism.
-As well, since we broadened our search criteria to include variations of neurodiver* such as ’neurodiversity’, ’neurodiverse’, and ’neurodivergent’, the search inevitably yielded posts discussing unrelated topics such as the NEURODIVER video game.
-The majority of these sentences were also published in 2023.
-7 Ethics Statement
-All the sentences in our dataset are publicly available, and we follow the methodologies of prior work in our data collection process  Atuhurra and Kamigaito (2024).
-We received IRB approval from our university’s review board and identified volunteer annotators through our association with various academic groups.
-Due to the sensitive nature of the content, we provided trigger warnings to our annotators and allowed them the freedom to annotate at their own pace.
-Additionally, we connected the members of each annotation team to enable discussions on the content and annotation process as needed.
-While our dataset and citations will be made available to the academic community, commercial use of the dataset is not allowed due to the size and nature of the data.
-As our knowledge of anti-autistic ableism continues to evolve, AUTALIC’s classification might become outdated.
-While we will be adding disclaimers if needed to reflect these changes, we still encourage researchers building upon our work to stay updated on the latest semantics by referring to the perspectives of autistic scholars, activists, and organizations.
+AuDHD A combination of autism and ADHD  Owens .
+Au, Âû Used by autistic individuals to self-identify as autistic  Union (2012).
+ND Neurodivergent, neurodiverse, or neurodiversity  Greally (2021)
+Table 5: Glossary of specialized terms used during annotation.
+A.5 Challenging Cases
+Here, we show examples of sentences with high disagreements among our annotators, along with their own notes on the aspects of each sentence that made it difficult to classify:
+Example 1
+“ABA can be very beneficial to low-functioning autists who can’t speak, use the bathroom, perform basic tasks, etc.”
+Promotes ABA and the “low-functioning” designation. Example 2
+“Autism can also lead to social issues like the autistic person saying the wrong thing at the wrong time with absolutely no utility involved.”
+Judgment that the autistic way is socially wrong.
+Example 3
+“Autism or ADHD or otherwise, you still have to learn basic etiquette.”
+Specific to cultural context: Implies NT (neurotypical) etiquette is required, but no expectation for NT to learn ways of relating that are natural to autistic people.
+Example 4
+“Basically, right after doing so much research, integrating with the autistic community, and accepting ASD as a part of myself, I was back to square one—left feeling like an idiot and immensely confused.”
+Slur against cognitive/intellectual disability, negativity associated with autistic identity, and medicalization of identity.
+A.6 Self-Agreement Scores
+In preliminary studies, we examined different labeling schemes for this task to identify the most efficient and effective annotation strategy.
+Our experiments revealed high levels of self-disagreement among annotators, as shown in Figure  5.
+The observed scores (M = -0.06, SD = 0.06) highlight the difficulty of the task and provide a meaningful baseline for comparison.
+Notably, our own annotation scores for Autalic were higher (M = 0.21, SD = 0.09), suggesting major improvement.
+Figure 5: The self-agreement scores among annotators in a preliminary study highlight the difficulty of this task.
+A.7 Annotation Platform
+Figure  6 shows an example of an annotation task on our platform with contextual sentences.
+Figure 6: An example of an annotation task on our platform containing the target sentence (green) and contextual sentences (white).

### v4

@@
 Abstract
-As our understanding of autism and ableism continues to increase, so does our understanding of ableist language towards autistic people.
+As our awareness of autism and ableism continues to increase, so does our understanding of ableist language towards autistic people.
 Such language poses a significant challenge in NLP research due to its subtle and context-dependent nature.
@@
 Autalic comprises 2,400 autism-related sentences collected from Reddit, accompanied by surrounding context, and annotated by trained experts with backgrounds in neurodiversity.
-Our comprehensive evaluation reveals that current language models, including state-of-the-art LLMs, struggle both to reliably identify anti-autistic ableism and to align with human judgments, underscoring their limitations in this domain.
-We publicly release Autalic along with the individual annotations.
-This dataset serves as a crucial step towards developing more inclusive and context-aware NLP systems that better reflect diverse perspectives.
+Our comprehensive evaluation reveals that current language models, including state-of-the-art LLMs, struggle to reliably identify anti-autistic ableism and diverge from human judgments, underscoring their limitations in this domain.
+We publicly release our dataset [^1] https://nrizvi.github.io/AUTALIC.html along with the individual annotations, providing an essential resource for developing more inclusive and context-aware NLP systems that better reflect diverse perspectives.
 Autalic: A Dataset for Anti-AUTistic Ableist Language In Context
@@
 Adding the surrounding sentences, as shown on the right, provides context revealing the original poster’s reference to the debunked vaccine-autism stereotype, which is tied to anti-autistic stigma Mann (2019); Davidson (2017).
-There are several critical frameworks used to define autism Lawson and Beckett (2021), including the medical model, which defines disability as a “disease” and is one of the most widely used in computer science research focusing on autism Rizvi et al.
-(2024); Spiel et al.
-(2019a); Sideraki and Drigas (2021); Anagnostopoulou et al.
+There are several critical frameworks used to define autism Lawson and Beckett (2021), including the medical model, which defines disability as a “disease” and remains one of the most widely used in computer science research focusing on autism Spiel et al.
+(2019a); Anagnostopoulou et al.
 (2020); Parsons et al.
+(2020); Sideraki and Drigas (2021); Sum et al.
 (2022); Williams et al.
-(2023); Sum et al.
+(2023); Rizvi et al.
 (2024).
 Since this framework defines autism as a deficit of skills, its applications in technology research largely focus on providing diagnosis and treatment to autistic people Baron-Cohen (1997); Begum et al.
-(2016); Rizvi et al.
 (2016); Spiel et al.
-(2019b).
+(2019b); Rizvi et al.
+(2024).
 This belief also posits neurotypical behaviors as the “norm” and autism as a “deficit” of these norms, thereby promoting neuronormativity instead of neurodiversity, which views all neurotypes as valid forms of human diversity Bottema-Beutel et al.
 (2021a); Walker (2014).
-To improve the alignment of AI research with neurodiversity, we present Autalic, a dataset of 2,400 autism-related sentences that we collect from various communities on Reddit, along with the original context of 2,014 immediately preceding and 2,400 following sentences.
-We aim to fill a critical gap in current NLP research, which has largely overlooked the nuanced and context-dependent nature of ableist speech targeting autistic individuals.
-Our dataset not only captures key contextual elements but also incorporates a comprehensive annotation process led by trained annotators with an understanding of the autistic community, ensuring higher reliability and relevance.
-Our final dataset contains all of the labels to capture the nuances in human perspectives, and allows Autalic to serve as a resource for researchers studying anti-autistic ableist speech, neurodiversity, or disagreements in general.
-Through a series of experiments with classical models and 4 LLMs, we find empirical evidence highlighting the difficulty of this task, and that LLMs are not reliable agents for such annotations.
-Our evaluations indicate that reasoning LLMs have the most consistent scores regardless of the language used in the prompt, thereby indicating a more thorough understanding of the different ways anti-autistic speech may be identified or may manifest in text.
+To move beyond the limitations of deficit-based models and better align NLP and AI research with neurodiversity, we present Autalic, a dataset of 2,400 autism-related sentences collected from Reddit, accompanied by 2,014 preceding and 2,400 following context sentences.
+The dataset addresses a critical gap in current NLP research, which has largely overlooked the nuanced and context-dependent nature of ableist speech targeting autistic individuals.
+Autalic captures key contextual elements and includes annotations by trained experts familiar with the autistic community, ensuring higher reliability and relevance.
+As Autalic contains all individual labels to capture the nuances in human perspectives, it can serve as a resource for researchers studying anti-autistic ableist speech, neurodiversity, or disagreements in general.
+Through a series of experiments with different baselines including 4 LLMs, our evaluation indicates that while some LLMs such as DeepSeek have the most consistent scores regardless of the language used in the prompt, thereby indicating a more thorough understanding of the different ways anti-autistic speech may be identified or may manifest in text, they remain unreliable agents for annotating such a task.
 We find that in-context learning examples provide mixed results in helping improve the task comprehension among LLMs.
+Our results provide empirical evidence highlighting the difficulty of this task.
 2 Related Work
@@
 It may include perpetuating stereotypes, using offensive language and slurs, or centering non-autistic people over the perspectives of autistic people Bottema-Beutel et al.
-(2021a); Rizvi et al.
-(2024); Darazsdi and Bialka (2023).
+(2021a); Darazsdi and Bialka (2023); Rizvi et al.
+(2024).
 While abusive language detection systems can help identify such speech, they are known to demonstrate bias Manerba and Tonelli (2021); Venkit et al.
@@
 (2023).
-Ableist language online varies, may manifest in different ways and is ever-evolving  Heung et al.
-(2024); Welch et al.
+Ableist language online varies, may manifest in different ways and is ever-evolving  Welch et al.
+(2023); Heung et al.
 (2024).
@@
 (2024).
-These limitations are also found in toxicity classifiers, which excel primarily at identifying explicit ableist speech but may otherwise perpetuate harmful social biases leading to content suppression Phutane et al.
+These limitations are also found in toxicity classifiers, which mainly identify explicit ableist speech but may otherwise perpetuate harmful social biases leading to content suppression Phutane et al.
 (2024).
@@
 This, unfortunately, can sometimes be due to a research design that overlooks intra-community and disabled people’s perspectives Mondal et al.
-(2022), as well as autistic people’s views, which may lead to harmful stereotypes Rizvi et al.
-(2024); Spiel et al.
-(2019b).
-We make a step towards addressing these issues by building a dataset that focuses on ableist speech and autism by including autistic people’s perspectives during the annotation processing as recommended by Davani et al.
+(2022), as well as autistic people’s views, which may lead to harmful stereotypes Spiel et al.
+(2019b); Rizvi et al.
 (2024).
-Autalic contains all its labels and will also be useful for researchers interested in leveraging disagreements for difficult classification tasks Leonardelli et al.
-(2021); Pavlick and Kwiatkowski (2019).
+We make a step towards addressing these issues by building a dataset that focuses on ableist speech and autism and including autistic people’s perspectives during the annotation process as recommended by Davani et al.
+(2023).
+Autalic contains all its labels and will also be useful for researchers interested in leveraging disagreements for difficult classification tasks Pavlick and Kwiatkowski (2019); Leonardelli et al.
+(2021).
 3 Autalic
@@
 (2024).
-The limitations of this method are detailed in Section  6.
 3.1.1 Data Collection Criteria
-We select Reddit as our source based on its popularity, focus on text-based content, and fewer API restrictions than X at the time of our data collection in January 2024.
-We search for keywords using the default search settings, which filters posts based on relevancy by prioritizing rare words in the search query, the age of the post, and the amount of likes and comments it has.
+We selected Reddit as our data source due to its popularity, emphasis on text-based content, and fewer API restrictions compared to X (formerly Twitter) at the time of data collection in January 2024.
+We performed keyword searches using Reddit’s default search settings, which prioritize relevance based on factors such as the rarity of terms in the query, the post’s age, and its number of upvotes and comment.
 [^2] https://support.reddithelp.com/hc/en-us/articles/19695706914196-What-filters-and-sorts-are-available The search terms include “autis*”, “ASD”, “aspergers”, and “disabilit*”; the full list is available in the Appendix, Table 4.
-Figure 2: An example of a sentence from our dataset. The search keyword is shown in red, while the word in blue is an example of a word defined in our glossary.
-We use the identified search terms to collect the target sentence instance containing our keywords, to be labeled by the annotators, and the sentences preceding or following target instances to provide additional context.
-We collect 2,400 target sentences, with 2,014 preceding and 2,400 following them.
-Finally, we split our dataset into three parts by randomly selecting and assigning 800 unique target sentences to create three segments that were each annotated by a group of three annotators.
-The average number of likes on each post included in the Autalic dataset is 1,611.59.
-Table 1 details the subreddits from which the most significant number of sentences were extracted from.
-With the exception of r/AmITheAsshole, all of the other subreddits are autism-related.
+Figure 2: An example sentence from our dataset. The search keyword is shown in red, while the word in blue is an example of a term defined in our glossary.
+We use the identified search terms to collect target sentence instances containing our keywords, which are then labeled by our annotators.
+To provide additional context, we also collect the sentences immediately preceding and following each target instance.
+In total, we collect 2,400 target sentences, along with 2,014 preceding and 2,400 following sentences.
+We then split the dataset into three parts by randomly selecting and assigning 800 unique target sentences to each of three annotation segments, with each segment labeled by a group of three annotators.
 SubReddit Sentence Count
@@
 Table 1: The subreddits with the most sentences included in the Autalic dataset and the number of sentences extracted from each.
+The average number of likes on each post included in the Autalic dataset is 1,611.59.
+Table 1 details the subreddits from which the most significant number of sentences were extracted from.
+With the exception of r/AmITheAsshole, all of the other subreddits are autism-related.
 3.1.2 Data Curation
@@
 We ensure that their involvement in our annotation process is voluntary, informed, and mutually beneficial.
-In particular, some participants choose to volunteer because they care deeply about the subject matter and wish to contribute to improving AI for autistic people.
+In particular, some participants chose to volunteer because they care deeply about the subject matter and wish to contribute to improving AI for autistic people.
 We select participants for whom this collaboration would provide relevant and valuable professional experience, grant them opportunities to engage in other aspects of the research, and provide mentorship and authorship recognition in accordance with the ACL guidelines.
@@
 We make them aware that they are free to withdraw from the study at any point.
-Our annotators are US-based, culturally diverse, and include people who grew up outside the United States.
+Our annotators are US-based, culturally diverse, and include people who grew up outside the US.
 They are all fluent in English.
@@
 Each annotator is assigned 800 unique sentences, with a goal of completing 200 annotations each week over four weeks.
-Annotators select from three possible labels for each sentence: “Ableist,” “Not Ableist,” or “Needs More Context.”
+Annotators select from three possible labels for each sentence: “Ableist”, “Not Ableist”, or “Needs More Context”.
 Ableist (1): We ask our annotators to select this label if a sentence contains ableist sentiments as defined by the Center for Disability Rights: “Ableism is a set of beliefs or practices that devalue and discriminate against people with physical, intellectual, or psychiatric disabilities and often rests on the assumption that disabled people need to be ‘fixed’ in one form or the other.” [^4] https://cdrnys.org/blog/uncategorized/ableism/
@@
 To calculate our agreement scores, we consolidated labels -1 and 0 together based on feedback from our annotators that unrelated sentences needing more context could be classified as not anti-autistic in a purely granular classification.
+While we use the majority label as the ground truth in our analysis, we release the individual labels from each annotator due to a growing interest in embracing disagreements for such classification tasks in NLP Plank et al.
+(2014); Pavlick and Kwiatkowski (2019); Leonardelli et al.
+(2021); Kralj Novak et al.
+(2022); Plank (2022).
 Label Definition Count
@@
 Table 2: An explanation of the labels used in our classification task and the resulting counts of each label from all 9 annotators combined.
-While we use the majority label as the ground truth in our analysis, in our public dataset, we will be releasing the individual labels from each annotator due to a growing interest in embracing disagreements for such classification tasks in NLP Leonardelli et al.
-(2021); Kralj Novak et al.
-(2022); Pavlick and Kwiatkowski (2019); Plank (2022); Plank et al.
-(2014).
 Our dataset contains 2,400 sentences labeled as containing anti-autistic ableist language or not.
 The labels are obtained by calculating the mode from the three annotators of each data segment.
-Using this methodology, 242 target sentences contain examples of anti-autistic ableist language (10% of total), and 2160 sentences do not (90% of total).
+Using this methodology, 242 target sentences contain examples of anti-autistic ableist language (10% of total), and 2,160 sentences do not (90% of total).
 3.2.4 Providing Context
-While we provide additional sentences for context, the annotators are instructed to annotate the target sentence exclusively and only refer to the other sentences for additional context, such as determining whether the sentence is part of an intra-community discussion or the use of figurative speech (i.e.
-sarcasm).
+While we provide additional sentences for context, the annotators are instructed to annotate the target sentence exclusively and only refer to the other sentences for additional context, such as determining whether the sentence is part of an intra-community discussion or the use of figurative speech (i.e., sarcasm).
 Figure 1 provides an example of a target sentence in context.
-In this example, it is difficult to determine whether or not the writer had ableist intent, as it can be interpreted in multiple ways.
-For example, they can be critiquing the medical model, as many autistic activists do, thereby making it non-ableist.
-Or they could be genuinely promoting ableist misrepresentations.
+That is, in this example, it is difficult to determine whether or not the writer had ableist intent, as it can be interpreted in multiple ways—they can be critiquing the medical model, as many autistic activists do, thereby making it non-ableist; or they could be genuinely promoting ableist misrepresentations.
 The contextual sentences help the annotators better understand the writer’s intent.
@@
   1. 1.
-    a tendency to use the medical model terminology or stereotypes (n=48)
+    a tendency to use the medical model terminology or stereotypes (in 48 cases),
   2. 2.
-    a need for additional context beyond the sentences we provided
+    a need for additional context beyond the sentences we provided.
 Figure 3: An example of a sentence from our dataset, shown in red, with a high level of disagreement among annotators.
@@
 Therefore, these sentences were ultimately classified as “not ableist” in Autalic.
-Our analysis reveals a moderately strong negative correlation between the task completion time and agreement with other annotators that is statistically significant ($R=-0.644$, $p$-value: $0.0096$).
+Our analysis reveals a moderately strong negative correlation between the task completion time and agreement with other annotators that is statistically significant (R=0.644R=-0.644, pp-value: $0.0096$).
 This highlights the importance of our orientation as we provided it simultaneously to the annotators.
 The annotators who completed their task immediately after our orientation had higher agreement.
-4 Experiments
-4.1 Experimental Setup
-We test the performance of different types of models on our dataset, including LLMs.
-These models are selected based on their diverse range of complexities.
-We also fine-tune BERT Devlin (2018) on our dataset to test classification.
-4.1.1 Models
-We use logistic regression (LR) with Bag of Words as the features, and using 80%-20% train-test split of the Autalic dataset.
+4 Experimental Setup
+We evaluate the performance of various baselines on our dataset using an 80%–20% train–test split. Additionally, we test four LLMs selected for their diverse range of model complexities.
+4.1 Baselines
+Logistic Regression (LR)
+We use logistic regression (LR) with Bag-of-Words (BoW) features for classification.
 Fine-tuned BERT
-We also utilize fine-tuned BERT Devlin et al.
-(2019) as a baseline for LLM experiment results, using an 80%-20% train-test split.
-The F1 scores for LR and BERT (both pretrained and fine-tuned) are presented in Table 3 alongside all LLM results.
-4.1.2 Prompting LLMs
-With our baseline established, we use the LLMs Gemma2  Team et al.
+We fine-tune BERT Devlin (2018) on our dataset to evaluate its performance on the classification task.
+4.2 Prompting LLMs
+We use four LLMs: Gemma2  Team et al.
 (2024), Mistral  Jiang et al.
@@
 (2024), and DeepSeek  Guo et al.
-(2025) to classify the sentences in our dataset, and we adjust the prompts to compare each LLM’s performance.
+(2025) to classify the sentences in our dataset.
+We adjust the prompts to compare each LLM’s performance.
 Due to limitations in computational resources, we are unable to fine-tune the LLMs.
@@
 (2023), we use three different kinds of prompts to measure the consistency LLMs have in their understanding of anti-autistic ableist speech.
-These prompts include person-first (i.e.
-’people with autism’), identity-first (i.e.
-’autistic people’), and conceptual (i.e.
-anti-autistic) language.
+These prompts include person-first (i.e., “people with autism”), identity-first (i.e., “autistic people”), and conceptual (i.e., “anti-autistic”) language.
 We keep the default parameters for each LLM to maintain consistency, and prompt them with the following questions:
@@
 In addition, we provide preceding and following context for each target sentence to the LLM to mimic the level of the information supplied to human annotators.
+Simple vs. ICL Prompting
 We run two sets of experiments with each LLM: one that uses zero-shot prompting, and another containing engineered prompts for in-context learning verbatim from the definitions and examples provided in our annotator orientation (Appendix Section A.2).
@@
 DeepSeek 0.55 0.56 0.55
-Table 3: The F1 scores of various models using person-first (PFL), identify-first (IFL), and conceptual anti-autistic (AA) prompts with and without in-context learning examples for each LLM.
+Table 3: The F1 scores of various models using person-first language (PFL), identity-first language (IFL), and conceptual anti-autistic (AA) prompts with and without in-context learning examples for each LLM.
 The best scores for each model are in bold.
 4.3 Experimental Results
-4.2.1 Fine-Tuning
-Our experiments reveal that utilizing BERT for this classification task can lead to high rates of censorship.
-As BERT (unlike the other LLMs tested) is not pre-trained with instructions, we obtain its results after fine-tuning on Autalic.
-While the pre-trained BERT model showed poor performance indicating that it was ineffective at identifying anti-autistic ableist speech, after fine-tuning on Autalic, the model’s performance improved dramatically across all metrics, indicating that it performs better at predicting ableist speech correctly, has fewer false positives, and has a higher sensitivity to recognizing ableist speech.
-4.2.2 Human-LLM Alignment
-Our assessment reveals that LLMs have low levels of alignment with human perspectives and the perspectives of other LLMs, which makes them unreliable agents for such classification tasks.
-We assess this alignment through a measurement using Cohen’s Kappa scores.
-The scores shown in Figure 4 indicate the highest level of alignment was demonstrated between Gemma2 and Mistral ($k=0.34$).
-No LLM demonstrated alignment with our human-annotated dataset, although DeepSeek’s alignment was notably higher than the others.
-Overall, the LLMs had low levels of agreement with human perspectives ($M=0.091,SD=0.110$).
-This indicates that LLMs with less than 10 billion parameters struggle with the task of classifying anti-autistic ableist language, even when provided with in-context examples.
+4.3.1 Fine-Tuning BERT
+The F1 scores for LR and BERT (both pretrained and fine-tuned) are presented in Table 3, alongside results from all LLMs.
+Our experiments reveal that using BERT for this classification task can initially result in high rates of censorship if deployed as-is.
+The pretrained BERT model performed poorly, indicating its ineffectiveness in identifying anti-autistic ableist speech.
+In contrast, fine-tuning BERT on Autalic led to significant performance improvements across all metrics, enabling more accurate predictions of ableist speech, fewer false positives, and greater sensitivity to subtle expressions of ableism.
+4.3.2 Humans vs. LLMs
+Our assessment reveals that LLMs show low levels of alignment with both human perspectives and the outputs of other LLMs, making them unreliable agents for this classification task if deployed as-is.
+We measure this alignment using Cohen’s Kappa scores.
+As shown in Figure 4, the highest level of agreement was observed between Gemma2 and Mistral ($\kappa=0.34$).
+No LLM demonstrated meaningful alignment with the human-annotated dataset, although DeepSeek’s agreement was notably higher than that of the other models.
+Overall, the LLMs exhibited low agreement with human judgments ($M=0.091$, $SD=0.110$).
+These results suggest that LLMs with fewer than 10 billion parameters struggle to classify anti-autistic ableist language, even when provided with in-context examples.
 4.3.3 In-Context Learning
-After providing the in-context learning examples, Llama3 (+22.96%) and Gemma2 (+12.68%) display the biggest relative improvement in F-1 scores, indicating that both models benefit from the examples.
-In particular, we find our ICL examples resulted in large relative improvements in consistency of scores regardless of the language used in the prompts for each LLM.
-For example, providing Llama with examples helped decrease its relative change in F1 scores from 67.49% to 17.4% when switching from PFL to conceptual prompts, indicating a better understanding of the connection between anti-autistic ableism and ableism toward autistic people.
-Figure 4: The mean Cohen’s Kappa scores of each LLM comparing the agreement with human annotators and other LLMs.
-4.2.4 Understanding Ableist and Anti-Autistic Speech
-Our results with prompt engineering reveal that anti-autistic ableism is too abstract of a concept for LLMs to recognize, providing empirical evidence that their current reasoning abilities are not inclusive of the perspectives of autistic people.
-LLMs struggle with identifying anti-autistic speech regardless of the terminology used, further indicating they are unreliable agents for data annotation tasks.
-Table 3 displays the results of prompt engineering using either person-first or identity-first language and using “anti-autistic” to describe this form of ableism in more conceptual terms.
-Notably, switching from PFL or IFL to conceptual language in the prompts resulted in the largest relative changes in scores.
-For instance, prompting Llama with "is this sentence anti-autistic?" instead of "is this sentence ableist toward people with autism?" resulted in a relative increase of 67.49%.
-These results show that LLMs struggle with understanding that “anti-autistic ableism" and “ableism toward autistic people" refer to the same phenomenon.
-Even after providing the ICL examples, the relative change in F1 scores between different prompts was as high as 32.66% for Gemma2.
-Interestingly, DeepSeek, the only reasoning LLM we test in our study, has the best results and highest consistency out of all the other LLMs.
+After providing in-context learning examples, Llama3 (+22.96%) and Gemma2 (+12.68%) show the largest relative improvements in F1 scores, indicating that both models benefit from the examples.
+In particular, our ICL examples also led to substantially more consistent scores across different prompt formulations for each LLM.
+For instance, supplying Llama3 with examples reduced its relative change in F1 score from 67.49% to 17.40% when switching from PFL to conceptual (AA) prompts, suggesting a better consideration of the connection between anti-autistic ableism and ableism toward autistic people.
+Figure 4: Mean Cohen’s Kappa scores for each LLM, comparing agreement with human annotators and other LLMs.
+4.3.4 Identifying Ableist and Anti-Autistic Speech
+Our results with prompt engineering reveal that anti-autistic ableism is too abstract a concept for LLMs to reliably recognize, providing empirical evidence that they do not adequately reflect the perspectives of autistic people.
+LLMs struggle to identify anti-autistic speech regardless of the terminology used, further indicating that they are unreliable agents for this data annotation task.
+Table 3 presents the results of prompt engineering using either person-first language (PFL), identity-first language (IFL), or the term “anti-autistic” (AA) to frame this form of ableism in more conceptual terms.
+Notably, switching from PFL or IFL to AA resulted in the largest relative changes in performance.
+For instance, prompting Llama with “Is this sentence anti-autistic?” instead of “Is this sentence ableist toward people with autism?” led to a relative increase of 67.49% in F1-score.
+These results indicate that LLMs struggle to frame anti-autistic ableism and ableism toward autistic people as referring to the same phenomenon.
+Even after introducing ICL examples, the relative change in F1-scores between different prompt types remained high—for example, 32.66% for Gemma2.
+Interestingly, DeepSeek has the best results and highest consistency out of all the others.
 Although its agreement with human annotators is low ($k:0.11$), it is still double that of all other LLMs, as shown in Figure  4.
 This highlights the difficulty of this task, as more advanced reasoning is required to understand the nuances of anti-autistic ableism, including the terminology we use to describe such speech.
+4.4 Error Analysis: Over-classification and Contextual Challenges for LLMs
+We analyze the top 10% of sentences in which our human annotators were in perfect agreement, but the LLMs disagreed with them.
+In doing so, we find many instances where the LLMs over-classified sentences as ableist.
+For example, Llama labeled 42 sentences as ableist, compared to 0 by our human annotators.
+Of these 42 sentences, 74-93% were also labeled as ableist by the other LLMs.
+Notably, 29 of these sentences involved intra-community discussions, highlighting the extent to which community voices may be censored if LLMs are used for content moderation.
+While two of the sentences did contain slurs, and 34 included other negatively connoted words—such as “burden”, “threat”, “harm”, or “stupid”—these terms were not used in an anti-autistic context.
+Figure 5 presents an example of such a sentence.
+Despite including negatively connoted terms, the sentence expresses an organization’s viewpoint that the author explicitly disagrees with.
+Since our annotators were instructed to consider context, they correctly labeled this sentence as not ableist.
+However, the LLMs misclassified it.
+This example demonstrates that even when given contextual information, LLMs tend to misclassify sentences based on the mere presence of negative words or connotations.
+This tendency contributes to their high rates of community censorship if deployed for moderation.
+Figure 5: An example of a sentence that LLMs classified as “anti-autistic”, in direct disagreement with all of our human annotators.
 4.5 Discussion
-Classifying anti-autistic ableist speech is challenging even within a Western, English-speaking context, since perceptions of what constitutes anti-autistic content differ significantly even among autistic individuals Keating et al.
+Classifying anti-autistic ableist speech is challenging even within a Western, English-speaking context, as perceptions of what constitutes anti-autistic content differ significantly—even among autistic individuals Keating et al.
 (2023).
-Factors such as personal experiences, cultural norms, and evolving discourse around autism advocacy can influence each individual’s perception of or sensitivity toward recognizing toxic speech, making it difficult to establish a consistent classification scheme Ousidhoum et al.
+Factors such as personal experience, cultural norms, and the evolving discourse around autism advocacy influence each individual’s perception of, and sensitivity to, toxic speech.
+This variability makes it difficult to establish a consistent classification scheme Kapp et al.
+(2013); Ousidhoum et al.
 (2019); Bottema-Beutel et al.
 (2021b); Taboas et al.
-(2023); Kapp et al.
 (2023).
-While developing Autalic, we standardize our definition of anti-autistic ableist speech by providing our annotators with an orientation and a glossary.
-These resources are developed in alignment with the perspectives of autistic people  Bottema-Beutel et al.
+While developing Autalic, we standardized our definition of anti-autistic ableist speech by providing annotators with an orientation and a glossary.
+These resources were created in alignment with the perspectives of autistic individuals Bottema-Beutel et al.
 (2021a); Taboas et al.
@@
 (2013).
-Our experiments demonstrate the importance of Autalic in aligning LLM performance to human expectations in the contexts of autism inclusion and ableist speech classification.
-Through our experiments, we provide empirical evidence of the current limitations of using LLMs and traditional classifiers to identify expressions of anti-autistic ableism.
-These limitations include: a misalignment with human perspectives, a lack of understanding of the concept of anti-autistic ableism, and a lack of agreement with each other even with in-context learning examples.
-Each of these limitations adds to the challenge of utilizing LLMs as reliable agents for such tasks.
-Standard pre-trained models such as showed poor performance, reinforcing the need for specialized fine-tuning.
-However, even after fine-tuning BERT on the Autalic dataset, our experiments reveal a high rate of false positives, which can lead to unfair censorship if BERT is employed for this task.
-Additionally, our results reveal that even state-of-the-art LLMs exhibit low agreement with human annotators on this task, further emphasizing the challenges of detecting subtle forms of ableism using generic models.
-DeepSeek has the best performance out of all the LLMs in our study, further demonstrating the difficulty of this task, as it is the only a reasoning-focused LLM in our study.
+Standard baselines such as logistic regression (LR) and BERT performed poorly, underscoring the need for domain-specific fine-tuning.
+Further, even after fine-tuning BERT on Autalic, our experiments reveal a high false positive rate, which could result in unfair censorship if used for content moderation.
+Furthermore, our results show that even state-of-the-art LLMs exhibit low agreement with human annotators, highlighting the difficulty of detecting subtle forms of ableism with general-purpose models.
+Among the models evaluated, DeepSeek achieved the best performance, further emphasizing the complexity of the task.
+These findings demonstrate the importance of Autalic in aligning LLM performance with human expectations in the context of autism inclusion and ableist speech classification.
+Our experiments provide empirical evidence of the current limitations of using LLMs and traditional classifiers to identify anti-autistic ableism.
+These limitations include misalignment with human perspectives, a lack of conceptual understanding of anti-autistic ableism, and inconsistency—even when provided with in-context learning examples.
+Each of these challenges hinders the reliability of LLMs for this task.
 5 Conclusion
-In this paper, we introduced Autalic, the first benchmark dataset focused specifically on the detection of anti-autistic ableist language in context.
-Through the collection and annotation of 2,400 sentences from Reddit, we aim to fill a critical gap in current NLP research and improve its alignment with neurodiversity.
-Looking forward, Autalic paves the way for significant advancements in content moderation systems, hate speech detection models, and research on ableism and neurodiversity.
-We envision this dataset as a cornerstone for future work in addressing bias against autistic individuals and fostering a more inclusive digital environment.
-By sharing this resource with the broader research community, we aim to catalyze the development of more equitable NLP systems that better serve underrepresented and marginalized groups.
+In this paper, we introduced Autalic, the first dataset specifically focused on anti-autistic ableist language in context.
+By collecting and annotating 2,400 Reddit sentences along with their surrounding context, we aim to fill a critical gap in current NLP research and improve its alignment with the principles of neurodiversity.
+By addressing bias against autistic individuals, Autalic marks a significant step forward in research on ableism within content moderation systems, hate speech detection models, and broader studies of neurodiversity and discrimination.
+We hope to support the development of more equitable NLP systems that better serve underrepresented and marginalized groups by sharing this resource with the research community,
 6 Limitations
-Despite doing our best to include a variety of ableist language against autistic people, our dataset can still show some selection bias Ousidhoum et al.
-(2020) as we relied on keywords and specific social media threads to collect our data.
-We recognize that, like other datasets presented at ACL, our work primarily reflects Western perspectives  Anderson-Chavarria (2022); Kirk et al.
-(2023); Thapa et al.
+Despite our efforts to include a variety of ableist language targeting autistic individuals, our dataset may still exhibit some selection bias Ousidhoum et al.
+(2020), as we relied on keyword searches and specific social media threads for data collection.
+We acknowledge that, like many datasets presented at ACL, our work primarily reflects Western perspectives Anderson-Chavarria (2022); Thapa et al.
 (2022); Aoyama et al.
-(2023); Takeshita et al.
-(2024); Park and Park (2023).
-Therefore, we do not claim that Autalic is generalizable across languages and cultures as anti-autistic and ableist speech may manifest and be perceived differently.
-However, we emphasize that Autalic is the first dataset to center the perspectives of autistic people in ableist speech detection, and move away from the tendency of AI to misclassify any text related to disability as ‘toxic’  Narayanan Venkit et al.
+(2023); Kirk et al.
+(2023); Park and Park (2023); Takeshita et al.
+(2024).
+Therefore, we do not claim that Autalic is generalizable across languages and cultures, as anti-autistic and ableist speech may manifest and be perceived differently in different contexts.
+However, we emphasize that Autalic is the first dataset to center the perspectives of autistic people in ableist speech detection, helping to move away from the tendency of AI systems to misclassify any disability-related text as “toxic” Hutchinson et al.
+(2020); Narayanan Venkit et al.
+(2023); Van Dorpe et al.
 (2023); Heung et al.
-(2024); Van Dorpe et al.
-(2023); Hutchinson et al.
 (2024).
-Therefore, it can serve as a reference for future work exploring ableist speech in diverse contexts such as other cultures and languages.
-Our dataset contains sentences from Reddit, and subreddits focusing on autism such as r/Aspergers, r/Autism, and r/AutismInWomen were among the most popular subreddits the posts were published in.
-Our search criteria included words like “r*tard” which is a slur broadly used against many intellectual disabilities and not specifically focused on autism.
-While we did our best to filter out irrelevant posts, since we broadened our search criteria to include variations of “neurodiver*” such as “neurodiversity”, “neurodiverse”, and “neurodivergent”, the search inevitably yielded posts discussing unrelated topics such as the NEURODIVER video game.
-The majority of these sentences were also published in 2023.
-While our dataset is small, the quality of our labels is high due to our standardized annotation protocols and resources that were developed using rigorous and iterative testing.
+As such, it can serve as a valuable reference for future work exploring ableist speech in diverse linguistic and cultural contexts.
+Our dataset consists of sentences sourced from Reddit, with many posts coming from autism-related subreddits such as r/Aspergers, r/Autism, and r/AutismInWomen.
+Our search criteria included terms like “r*tard”, a slur broadly used against people with intellectual disabilities, not exclusively targeting autism.
+While we made efforts to filter out irrelevant posts, our inclusion of broader search terms like “neurodiver” (e.g., “neurodiversity”, “neurodiverse”, and “neurodivergent”) also returned unrelated content, such as discussions about the NEURODIVER video game.
+Most of the dataset’s content was published in 2023.
+Although our dataset is relatively small, the quality of the annotations is high, due to standardized protocols and resources developed through rigorous, iterative testing.
 7 Ethics Statement
-The data collected for our small-scale and non-commerical research is in compliance with Reddit’s API limits and policies  Reddit (2023a, b).
-All the sentences in our dataset are publicly available, and we follow the methodologies of prior work in our data collection process  Atuhurra and Kamigaito (2024).
-We specifically recruit annotators for whom this collaboration and resulting paper authorship would be mutually beneficial, and provide a comprehensive overview of the task to ensure they make an informed decision to participate.
-Some of our annotators chose to volunteer as they care deeply about neurodiversity and autism inclusion.
-We received IRB approval from our university’s review board and identified volunteer annotators through our association with various academic groups.
-Given the sensitivity of the content, we provided annotators with appropriate trigger warnings, ensuring they could work at their own pace or withdraw from the study if necessary.
-Additionally, we connected the members of each annotation team to enable discussions on the content and annotation process as needed.
-While our dataset and citations will be made available to the academic community, commercial use of the dataset is not allowed due to the size and nature of the data.
-As our knowledge of anti-autistic ableism continues to evolve, Autalic’s classification might become outdated.
-While we will be adding disclaimers if needed to reflect these changes, we still encourage researchers building upon our work to stay updated on the latest semantics by referring to the perspectives of autistic scholars, activists, and organizations.
-Refer to our Appendix for our Guidelines for Responsible Use.
+The data collected for our small-scale, non-commercial research complies with Reddit’s API limits and policies Reddit (2023a, b).
+All sentences in our dataset are publicly available, and our data collection process follows the methodologies established by prior work Atuhurra and Kamigaito (2024).
+We obtained IRB approval from our university’s ethics review board and recruited volunteer annotators through our connections with various academic groups.
+Given the sensitivity of the content, we provided annotators with appropriate trigger warnings and allowed them to work at their own pace or withdraw from the study at any time.
+Additionally, we facilitated communication within each annotation team to support discussion of the content and the annotation process as needed.
+We specifically recruited annotators for whom this collaboration—and potential paper authorship—would be mutually beneficial, and we provided a comprehensive overview of the task to ensure informed consent.
+Some annotators chose to volunteer due to their strong commitment to neurodiversity and autism inclusion.
+While our dataset and associated citations will be made available to the academic community, commercial use—including use by state actors in high-risk domains (e.g., healthcare)—is strictly prohibited.
+Furthermore, as our understanding of anti-autistic ableism continues to evolve, the classifications in Autalic may become outdated.
+We will include disclaimers to reflect such developments as needed, but we strongly encourage researchers building on our work to stay informed by consulting current perspectives from autistic scholars, activists, and organizations.
+Please refer to the Appendix for our Guidelines for Responsible Use.
 Appendix A Appendix
@@
   • The impact of the project, and its target end-users
-Contact the First Author: All requests and inquiries should be directed to the first author, as listed in the dataset documentation or project website, available here: [REDACTED FOR PRIVACY]
+Contact the First Author: All requests and inquiries should be directed to the first author, as listed in the dataset documentation or project website, available here  [^5] https://nrizvi.github.io/AUTALIC.html.
 A.1.8 Liability and Disclaimer
@@
 In preliminary studies, we examined different labeling schemes for this task to identify the most efficient and effective annotation strategy.
-Our experiments revealed high levels of self-disagreement among annotators, as shown in Figure  5.
+Our experiments revealed high levels of self-disagreement among annotators, as shown in Figure  6.
 The observed scores (M = -0.06, SD = 0.06) highlight the difficulty of the task and provide a meaningful baseline for comparison.
@@
 A.7 Annotation Platform
-Figure  6 shows an example of an annotation task on our platform with contextual sentences.
+Figure  7 shows an example of an annotation task on our platform with contextual sentences.
 Figure 7: An example of an annotation task on our platform containing the target sentence (green) and contextual sentences (white).